### v1

# The uniqueness theorem for Kasparov theory

Abstract.
Answering a question of Carrión et al in their recent landmark paper on $\mathrm{C}^{*}$-algebra classification, we prove a general uniqueness theorem for $KK$-theory. Given arbitrary separable $\mathrm{C}^{*}$-algebras $A$ and $B$ and a Cuntz pair consisting of two absorbing representations $\varphi,\psi:A\to\mathcal{M}(B\otimes\mathcal{K})$, the induced element of $KK(A,B)$ vanishes if and only if $\varphi$ and $\psi$ are strongly asymptotically unitarily equivalent. This improves upon the Lin–Dadarlat–Eilers stable uniqueness theorem. The conclusion is deduced by first showing the $K_{1}$-injectivity of an auxiliary $\mathrm{C}^{*}$-algebra associated to the $\mathrm{C}^{*}$-pair $(A,B)$, which is sometimes called the Paschke dual algebra in the literature. Most of the article is concerned with the treatment of an umbrella theorem, which yields such a uniqueness theorem for other variants of $KK$-theory. This encompasses nuclear $KK$-theory, ideal-related $KK$-theory, equivariant $KK$-theory, or any combinations thereof.
2020 Mathematics Subject Classification:
19K35, 46L55, 46L35
Funded by the European Union. Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the EU nor the ERC can be held responsible for them.
Introduction
Kasparov theory, also known as $KK$-theory, is a very powerful invariant for $\mathrm{C}^{*}$-algebras. It is a bivariant version of topological $K$-theory, which includes ordinary $K$-theory and $K$-homology, alongside with Brown–Douglas–Fillmore theory [5], as special cases. Invented by Kasparov [30, 31], it can be viewed as either a bifunctor from (separable) C*-algebras to abelian groups, or as a way to equip the class of (separable) $\mathrm{C}^{*}$-algebras with a more flexible notion of homomorphisms. In fact, given a second-countable locally compact group $G$, Kasparov’s construction also yields such a bifunctor on the category of all (separable) $G$-$\mathrm{C}^{*}$-algebras, which provided impactful applications on classical problems such as the Novikov conjecture. The original approach by Kasparov, commonly referred to as the Kasparov picture, is still the most common way of treating this subject, but there are other equivalent appproaches that are geared towards different types of applications.
In this article we are mainly interested in the Cuntz picture [8] and the Cuntz–Thomsen picture [42] for (equivariant) $KK$-groups. In Cuntz’ framework, elements in Kasparov’s group $KK(A,B)$ for separable $\mathrm{C}^{*}$-algebras $A$ and $B$ can be described as homotopy classes of so-called Cuntz pairs, i.e., pairs of *-homomorphisms $\varphi,\psi:A\to\mathcal{M}(B\otimes\mathcal{K})$ such that $\varphi(a)-\psi(a)\in B\otimes\mathcal{K}$ for all $a\in A$. For equivariant $KK$-theory, one considers pairs of cocycle representations satisfying a similar condition; see also [41, Section 6]. A groundbreaking theorem concerning the analytical properties of $KK$-theory is the so-called stable uniqueness theorem proved independently by Lin [34] and Dadarlat–Eilers [12, 13]. It states that if a Cuntz pair $(\varphi,\psi)$ represents the zero element in $KK(A,B)$, then $\varphi$ and $\psi$ are homotopic via conjugating one of the maps with a unitary path in $\mathbf{1}+B\otimes\mathcal{K}$ after stabilizing both $\varphi$ and $\psi$ with a so-called absorbing representation. Note that an elegant treatment of the stable uniqueness theorem was recently included in [10]. Over the years, this theorem has become one of the central pieces of methodology in the structure and classification theory of nuclear $\mathrm{C}^{*}$-algebras; see for example [46, 26, 25, 16, 14, 22, 38, 23, 18, 24, 6]. More recently, I have generalized the stable uniqueness theorem to the setting of $\mathrm{C}^{*}$-dynamics together with Gabe in [21], which was one of the key conceptual ingredients behind our dynamical Kirchberg–Phillips theorem [20]. Further interesting applications of this theorem may be expected to come in the realm of $\mathrm{C}^{*}$-dynamics and their classification. Its counterpart for actions of unitary tensor categories followed soon afterwards [33] and may become instrumental to classify unitary tensor category actions on Kirchberg algebras.
When one considers the stable uniqueness theorem (either in its $\mathrm{C}^{*}$-algebraic or $\mathrm{C}^{*}$-dynamical form), the role of the absorbing representation is that of a necessary largeness criterion for the representations in question. The notion of an absorbing representation — recalled after subsection 1.1 for $\mathrm{C}^{*}$-dynamics — is at the same time inspired by the conclusion of Voiculescu’s fundamental theorem [47]. It can be interpreted as asserting that every unitizably full [^1] In this specific context, this can be interpreted as one of two more familiar statements. If the $\mathrm{C}^{*}$-algebra is non-unital, then the representation needs to be faithful and essential. If the $\mathrm{C}^{*}$-algebra is unital, then it also needs to have the property that the representation is supported on a subspace with infinite-dimensional orthogonal complement. representation of a separable $\mathrm{C}^{*}$-algebra on the Hilbert space $\ell^{2}(\mathbb{N})$ is absorbing. It is known that all elements of a given $KK$-group can be represented by a Cuntz pair of absorbing representations. When one is faced with applications to the classification of nuclear $\mathrm{C}^{*}$-algebras, such an absorbing behavior can be detected via Elliott–Kucerovsky’s “purely largeness” criterion [15] and its known variants; see also [3, Subsection 4.2]. For the purpose of this article, one of the primary applications to have in mind in classification is the approach of Carrión et al [6], where part of the technical work is about getting closer to a setup where one can compare a pair of naturally occurring absorbing representations via $KK$-theory. The stable uniqueness theorem allows one to obtain a strong form of equivalence for the pair of representations, but in general only after first stabilizing with an auxiliary representation. The process to “destabilize”, so to speak, may require a substantial amount of extra work involving the specific assumptions related to the application. Schafhauser demonstrated in [38] that this issue can be circumvented under a UHF-stability assumption. Building on this idea, it is shown in [6, Theorem 5.15] that one can omit the stabilization with the auxiliary representation at the expense of adding a tensor copy of the Jiang–Su algebra $\mathcal{Z}$ to the codomain. Since said article concerns the classification of $\mathrm{C}^{*}$-algebras that are ultimately Jiang–Su stable, this extra tensor copy of $\mathcal{Z}$ can be removed in a later phase of the application with some extra steps. However, it is a natural question to what extent it is really necessary to stabilize with an auxiliary representation when one compares a pair of representations that are absorbing to begin with. This was asked as [6, Question 5.17] and [39, Problem LXII]. One would refer to the desired statement below as a “$KK$-uniqueness theorem”.
Question A.
Suppose that $A$ and $B$ are separable $\mathrm{C}^{*}$-algebras. Let $\varphi,\psi:A\to\mathcal{M}(B\otimes\mathcal{K})$ be two absorbing representations forming a Cuntz pair. If $\varphi,\psi]=0$ in $KK(A,B)$, is it true that $\varphi$ and $\psi$ are strongly asymptotically unitarily equivalent?
The discussions in the aforementioned references point towards the main technical obstacle being the $K_{1}$-injectivity of the auxiliary $\mathrm{C}^{*}$-algebra
$\mathcal{Q}(B\otimes\mathcal{K})\cap\bar{\varphi}(A)^{\prime},$
which had been a problem considered beforehand in other related works [37, 19, 35, 36] (here $\mathcal{Q}$” denotes the corona algebra and $\bar{\varphi}$ is the induced homomorphism). More recently, I demonstrated with Farah in [17, Theorem 5.5] that under the assumption of Jiang–Su stability on $B$, this auxiliary object is separably Jiang–Su stable, which implies $K_{1}$-injectivity (by [29] or [6, Theorem 4.8]) and thus the $KK$-uniqueness theorem is indeed true when the codomain algebra $B$ is Jiang–Su stable.
The first and main goal of this article is to answer Introduction in the affirmative, by way of showing that the technical obstacle related to $K_{1}$-injectivity has a positive solution for arbitrary pairs of separable $\mathrm{C}^{*}$-algebras. We mention briefly that as per the discussion around [6, Question 5.17], Introduction has a matching version involving a similar statement for $KL$-theory, where the desired conclusion is instead a version of approximate unitary equivalence of the two involved presentations. While this kind of $KL$-uniqueness is also very useful in applications, it is implicit from that discussion that it can be deduced as the consequence of a general answer to Introduction. This is the reason why $KL$-theory plays no explicit role in the rest of this article. To describe the approach to answer Introduction, it is more useful to consider the $\mathrm{C}^{*}$-algebra $\mathscr{D}_{\varphi}\subseteq\mathcal{M}(B\otimes\mathcal{K})$ of all multipliers whose induced class belongs to $\mathcal{Q}(B\otimes\mathcal{K})\cap\bar{\varphi}(A)^{\prime}$; see section 2. The main result of the second section is the following consequence of section 2 and section 2.
Theorem B.
Let $A$ and $B$ be separable $\mathrm{C}^{*}$-algebras. If $\varphi:A\to\mathcal{M}(B\otimes\mathcal{K})$ is (unitally) absorbing, then $\mathscr{D}_{\varphi}$ and all its quotients are $K_{1}$-injective.
Combined with the detailed discussions appearing after [6, Question 5.17] and [39, Problem LXII], this implies a positive answer to Introduction. The added layers of generality within section 2 include other recent results of this type like [35, Theorems 2.5, 2.9] and [36, Theorem 3.28] as special cases.
Before we move on to the remaining content of the article, I shall say a few words about the approach to prove Introduction. The core argument behind the proof draws heavy inspiration from the proof of Cuntz–Higson’s theorem [11] that the unitary group of every stable multiplier algebra $\mathcal{M}(B\otimes\mathcal{K})$ is connected in the norm topology. If one analyzes said proof carefully, one can interpret it as a two-step approach: The first (and relatively easy) step is to observe that the $K$-groups of $\mathcal{M}(B\otimes\mathcal{K})$ vanish, which is a consequence of the fact that it possesses an infinite repeat endomorphism, which automatically trivializes the identity map in $K$-theory. The second (and more difficult) step is to show that $\mathcal{M}(B\otimes\mathcal{K})$ is $K_{1}$-injective. Since it is properly infinite as a unital $\mathrm{C}^{*}$-algebra, there are known criteria to verify $K_{1}$-injectivity (these are nicely summarized in [2]), one of which was the original criterion verified by Cuntz–Higson. In essence, the proof of Introduction (say for the $K_{1}$-injectivity of the $\mathrm{C}^{*}$-algebra $\mathscr{D}_{\varphi}$ itself) can be seen to mimic Cuntz–Higson’s argument, with the added point that we need to keep track of the relative commutant condition appearing in the definition of $\mathscr{D}_{\varphi}$. Fortunately, this works in remarkable generality as long as $\varphi$ is sufficiently similar to an infinite repeat representation, which is automatically the case when $\varphi$ is absorbing. With this requisite in hand, the main difference between Cuntz–Higson’s original argument and the one appearing here is that instead of utilizing arbitrary increasing approximate units in $B\otimes\mathcal{K}$, we utilize approximate units that are quasicentral relative to $\varphi(A)$. Since the proof involves no further assumptions, the conclusion about $K_{1}$-injectivity can be drawn in a much more flexible situation (described in section 2), which is applicable to other auxiliary $\mathrm{C}^{*}$-algebras that play an analogous role for other types of $KK$-theory such as nuclear, ideal-related, or equivariant $KK$-theory. The argument does however very much depend on the fact that the $\mathrm{C}^{*}$-algebra in question comes from a stable multiplier algebra. Hence I would like to mention, for those readers wondering, that the proof behind Introduction likely sheds no light on the general $K_{1}$-injectivity problem for properly infinite $\mathrm{C}^{*}$-algebras; see [2, Question 2.9] and [39, Problem LXI].
In the rest of the article, I aim to treat the framework concerning the $KK$-uniqueness theorem in the most general setting possible for $\mathrm{C}^{*}$-algebras or $\mathrm{C}^{*}$-dynamics. Having in mind other known variants of the stable uniqueness theorem and their applications, it can be just as useful to look at $KK$-groups with extra structure, for instance Skandalis’ nuclear $KK$-theory [40] or Kirchberg’s ideal-related $KK$-theory [32]; see [18, Section 13] for a unified treatment. It stands to reason that in the long term, such $KK$-groups with enriched structure will also play a similar role in the structure and classification of $\mathrm{C}^{*}$-dynamics. The third section introduces an umbrella framework of sorts, which to my knowledge unifies all known variants of (equivariant) $KK$-theory that have received significant attention in the literature. Given two separable $G$-$\mathrm{C}^{*}$-algebras $(A,\alpha)$ and $(B,\beta)$ with $\beta$ strongly stable, the needed input data consists of some set $\mathfrak{C}$ of cocycle representations $(\varphi,\mathbbmss{u}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ containing all elements of the form $(0,\mathbbmss{x})$, and which is downward closed under weak containment (subsection 1.1). A set $\mathfrak{C}$ satisfying these abstract properties is referred to as an eligible set. The modified $KK$-group $KK^{G}(\mathfrak{C};\alpha,\beta)$ is then built, in analogy to [42], from Cuntz pairs of cocycle representations in $\mathfrak{C}$, with a homotopy relation that requires the continuous paths to consist of cocycle representations in $\mathfrak{C}$ (called $\mathfrak{C}$-homotopy). As is discussed further in section 3, this general construction has as special cases nuclear and ideal-related $KK$-theory. In the third section it is shown that one of Kasparov’s fundamental observations extends to this variant of $KK$-theory, namely that $\mathfrak{C}$-homotopy is equivalent to operator homotopy after stabilizing with a representation in $\mathfrak{C}$. This generalizes the results in [21, Section 5] and is based on just slightly adjusting the approach therein to leverage the properties of an eligible set.
As a consequence of results in [21], as long as the involved $\mathrm{C}^{*}$-algebras are separable, every eligible set has some absorbing element, i.e., a representation $(\varphi,\mathbbmss{u})\in\mathfrak{C}$ that absorbs every representation in $\mathfrak{C}$. Just as before, it can be shown that every element of $KK^{G}(\mathfrak{C};\alpha,\beta)$ has a representative Cuntz pair consisting of two absorbing elements in $\mathfrak{C}$. In the fourth section, the results obtained so far are combined with a few further analytic arguments to prove the appropriate $KK$-uniqueness theorem for these generalized $KK$-groups. The following is the separable version of section 4.
Theorem C.
Let $G$ be a second-countable locally compact group. Let $\alpha:G\curvearrowright A$ and $\beta:G\curvearrowright B$ be two actions on separable $\mathrm{C}^{*}$-algebras, and suppose that $\beta$ strongly stable. Let $\mathfrak{C}$ be an eligible set of cocycle representations $(A,\alpha)\to(\mathcal{M}(B),\beta)$. Let $(\varphi,\mathbbmss{u})$ and $(\psi,\mathbbmss{v})$ be two absorbing elements of $\mathfrak{C}$ that form an anchored $(\mathfrak{C};\alpha,\beta)$-Cuntz pair. Then $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v})]=0$ in $KK^{G}(\mathfrak{C};\alpha,\beta)$ if and only if $(\varphi,\mathbbmss{u})$ and $(\psi,\mathbbmss{v})$ are strongly asymptotically unitarily equivalent.
[…]
As an outlook for future applications, Hua–White’s recent article [28] serves as an excellent template. Therein, the authors prove a special case of the $KK$-uniqueness theorem, tailored to their intended application to classify certain nuclear embeddings of exact $\mathrm{C}^{*}$-algebras into ultrapowers of matrix algebras. Their setup did not match the assumptions required for the classification machinery in [6], but their work introduces various novel ideas that demonstrate how the scope of such classification results can be extended if one has access to the general $KK$-uniqueness theorem. I thus hope that the results of this article will be used in future works to obtain a more general classification of nuclear embeddings of exact $\mathrm{C}^{*}$-algebras into ultraproducts of $\mathrm{C}^{*}$-algebras that need not be Jiang–Su stable.
1. Preliminaries
Notation 1.1.
Throughout, $G$ will denote a second-countable, locally compact group unless specified otherwise. Normal capital letters like $A,B,C$ denote $\mathrm{C}^{*}$-algebras that are frequently assumed to be separable. The multiplier algebra of $A$ is denoted as $\mathcal{M}(A)$. We write $\mathcal{U}(\mathbf{1}+A)$ for the set of all unitaries in the proper unitization $A^{\dagger}$ whose scalar part is $\mathbf{1}$, which can be canonically identified with the unitary group of $A$ if it was already unital. If we are given a (closed two-sided) ideal $J\triangleleft A$ and two elements $a,b\in A$, then the expression $a\equiv_{J}b$ is short-hand for $a-b\in J$. The symbol $\mathcal{K}$ denotes the $\mathrm{C}^{*}$-algebra of compact operators on a separable infinite-dimensional Hilbert space $\mathcal{H}$, and we write $\mathcal{K}(\mathcal{H})$ when a specific description of $\mathcal{H}$ is relevant in a given context. A $\mathrm{C}^{*}$-algebra $A$ is called stable when $A\cong A\otimes\mathcal{K}$. Greek letters such as $\alpha,\beta,\gamma$ are used for point-norm continuous maps $G\to\operatorname{Aut}(A)$, most often group actions, which uniquely induces maps into $\operatorname{Aut}(\mathcal{M}(A))$ as well. Depending on the situation, we may denote $\operatorname{id}_{A}$ either for the identity map on $A$ or the trivial $G$-action on $A$. We denote by $A^{\alpha}$ or $\mathcal{M}(A)^{\alpha}$ the $\mathrm{C}^{*}$-subalgebra of fixed points (in $A$ or $\mathcal{M}(A)$) with respect to an action $\alpha$. Normal alphabetical letters such as $u,v,U,V$ are often used for unitary elements in some $\mathrm{C}^{*}$-algebra $A$. If either $u\in\mathcal{U}(\mathcal{M}(A))$ or $u\in\mathcal{U}(\mathbf{1}+A)$, we denote by $\operatorname{Ad}(u)$ the induced inner automorphism of $A$ given by $a\mapsto uau^{*}$. Double-struck letters such as $\mathbbmss{u},\mathbbmss{v},\mathbb{U},\mathbb{V}$ are used for strictly continuous maps $G\to\mathcal{U}(\mathcal{M}(A))$, most often cocycles with respect to an action $\alpha:G\curvearrowright A$, i.e., one has the cocycle identity $\mathbbmss{u}_{gh}=\mathbbmss{u}_{g}\alpha_{g}(\mathbbmss{u}_{h})$ for all $g,h\in G$. Under this assumption, one obtains a new (cocycle perturbed) action $\alpha^{\mathbbmss{u}}:G\curvearrowright A$ via $\alpha^{\mathbbmss{u}}_{g}=\operatorname{Ad}(\mathbbmss{u}_{g})\circ\alpha_{g}$.
Although we will introduce equivariant $KK$-theory in this section via the Cuntz–Thomsen picture [8, 9, 27, 42], we will implicitly assume that the reader has some existing passing familiarity with it; see for example [1]. We will assume that the reader is at least fluent in the language of $K$-theory for $\mathrm{C}^{*}$-algebras.
We will make use of the following fundamental result due to Kasparov concerning quasicentral approximate units coming from ideals invariant under a group action.
Lemma 1.2 (see [31, Lemma 1.4] and its proof).
Let $\beta:G\curvearrowright B$ be an action on a $\sigma$-unital $\mathrm{C}^{*}$-algebra and let $e_{n}\in B$ be any countable increasing approximate unit. Then for any separable $\mathrm{C}^{*}$-subalgebra $D\subseteq\mathcal{M}(B)$, there exists a countable increasing approximate unit of positive contractions $h_{n}\in B$ belonging to the convex hull of $\{e_{n}\mid n\geq 1\}$ satisfying
$\lim_{n\to\infty}\|[h_{n},d]\|=0\quad\text{and}\quad\lim_{n\to\infty}\max_{g\in K}\|h_{n}-\beta_{g}(h_{n})\|=0$
for all $d\in D$ and compact sets $K\subseteq G$. If one has $e_{n+1}e_{n}=e_{n}$ for all $n\geq 1$, then one can arrange $h_{n+1}h_{n}=h_{n}$ for all $n\geq 1$.
1.1. Some cocycle representation theory
Definition 1.3 (see [41, Section 1]).
Let $\alpha:G\curvearrowright A$ and $\beta:G\curvearrowright B$ be two actions on $\mathrm{C}^{*}$-algebras. A cocycle representation $(\varphi,\mathbbmss{u}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ consists of a *-homomorphism $\varphi:A\to\mathcal{M}(B)$ and a strictly continuous $\beta$-cocycle $\mathbbmss{u}:G\to\mathcal{U}(\mathcal{M}(B))$ satisfying $\operatorname{Ad}(\mathbbmss{u}_{g})\circ\beta_{g}\circ\varphi=\varphi\circ\alpha_{g}$ for all $g\in G$. For any unitary $W\in\mathcal{U}(\mathcal{M}(B))$, we can obtain a new cocycle representation defined by conjugation as
$\operatorname{Ad}(W)\circ(\varphi,\mathbbmss{u}):=(\operatorname{Ad}(W)\circ\varphi,\{W\mathbbmss{u}_{g}\beta_{g}(W)^{*}\}_{g\in G}).$
If a given cocycle representation can be obtained from another via this procedure, then they are called unitarily equivalent. If there exists a cocycle representation $(\varphi,\mathbbmss{u})$ such that $\varphi$ is an isomorphism from $A$ to $B$, then $\alpha$ and $\beta$ are called cocycle conjugate.
Notation 1.4.
We say that an action $\beta:G\curvearrowright B$ is strongly stable if $(B,\beta)$ is conjugate (i.e., equivariantly isomorphic) to $(B\otimes\mathcal{K},\beta\otimes\operatorname{id}_{\mathcal{K}})$.
An action $\beta:G\curvearrowright B$ is strongly stable if and only if there is a sequence of isometries $r_{n}\in\mathcal{M}(B)^{\beta}$ such that $\mathbf{1}=\sum_{n=1}^{\infty}r_{n}r_{n}^{*}$ holds in the strict topology; see [21, Remark 2.4]. Furthermore, it was observed in [20, Proposition 1.4] that every $G$-action on a stable $\mathrm{C}^{*}$-algebra is cocycle conjugate to a strongly stable action.
Definition 1.5 (cf. [42, Lemma 3.4]).
Let $\beta:G\curvearrowright B$ be an action on a $\mathrm{C}^{*}$-algebra. Suppose that there exists a unital inclusion $\mathcal{O}_{2}\subseteq\mathcal{M}(B)^{\beta}$. For two isometries $t_{1},t_{2}\in\mathcal{M}(B)^{\beta}$ with $t_{1}t_{1}^{*}+t_{2}t_{2}^{*}=\mathbf{1}$, we may consider the $\beta$-equivariant *-homomorphism
$\mathcal{M}(B)\oplus\mathcal{M}(B)\to\mathcal{M}(B),\quad b_{1}\oplus b_{2}\mapsto b_{1}\oplus_{t_{1},t_{2}}b_{2}:=t_{1}b_{1}t_{1}^{*}+t_{2}b_{2}t_{2}^{*}.$
Up to unitary equivalence with a unitary in $\mathcal{M}(B)^{\beta}$, this *-homomorphism does not depend on the choice of $t_{1}$ and $t_{2}$. [^2] If $v_{1},v_{2}\in\mathcal{M}(B)^{\beta}$ are two other isometries with $v_{1}v_{1}^{*}+v_{2}v_{2}^{*}=\mathbf{1}$, then the unitary equivalence between $\oplus_{t_{1},t_{2}}$” and $\oplus_{v_{1},v_{2}}$” is implemented by $w=t_{1}v_{1}^{*}+t_{2}v_{2}^{*}\in\mathcal{M}(B)^{\beta}$. One refers to the element $b_{1}\oplus_{t_{1},t_{2}}b_{2}$ as the Cuntz sum of the two elements $b_{1}$ and $b_{2}$ (with respect to $t_{1}$ and $t_{2}$). Now let $\alpha:G\curvearrowright A$ be another action on a $\mathrm{C}^{*}$-algebra, and $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ two cocycle representations. We likewise define the (pointwise) Cuntz sum
$(\varphi,\mathbbmss{u})\oplus_{t_{1},t_{2}}(\psi,\mathbbmss{v})=(\varphi\oplus_{t_{1},t_{2}}\psi,\mathbbmss{u}\oplus_{t_{1},t_{2}}\mathbbmss{v}):(A,\alpha)\to(\mathcal{M}(B),\beta),$
which is easily seen to be another cocycle representation. Since its unitary equivalence class does not depend on the choice of $t_{1}$ and $t_{2}$, we will often omit $t_{1}$ and $t_{2}$ from the notation if it is clear from context that a given statement is invariant under unitary equivalence.
Similarly to how we formed Cuntz sums of two elements, we may also form countably infinite sums by a similar method if the underlying action is strongly stable.
We shall, from this point and throughout the article, always assume that $\beta:G\curvearrowright B$ denotes a strongly stable action on a $\sigma$-unital $\mathrm{C}^{*}$-algebra and that $\alpha:G\curvearrowright A$ is an action on a separable $\mathrm{C}^{*}$-algebra, unless specified otherwise.
Definition 1.6.
Let $t_{n}\in\mathcal{M}(B)^{\beta}$ be any sequence of isometries such that $\sum_{n=1}^{\infty}t_{n}t_{n}^{*}=\mathbf{1}$ in the strict topology. Then we have a $\beta$-equivariant *-homomorphism
$\ell^{\infty}(\mathbb{N},\mathcal{M}(B))\to\mathcal{M}(B),\quad(b_{n})_{n\geq 1}\mapsto\sum_{n=1}^{\infty}t_{n}b_{n}t_{n}^{*},$
which does not depend on the choice of $t_{n}$ up to unitary equivalence with a unitary in $\mathcal{M}(B)^{\beta}$. [^3] Similarly as before, if $v_{n}\in\mathcal{M}(B)^{\beta}$ is another sequence of isometries satisfying the same relation, then the unitary $w=\sum_{n=1}^{\infty}t_{n}v_{n}^{*}$ implements this equivalence. For any sequence of cocycle representations $(\varphi^{(n)},\mathbbmss{u}^{(n)}):(A,\alpha)\to(\mathcal{M}(B),\beta)$, we may hence define the countable sum
$(\Phi,\mathbb{U})=\bigoplus_{n=1}^{\infty}(\varphi^{(n)},\mathbbmss{u}^{(n)}):(A,\alpha)\to(\mathcal{M}(B),\beta)$
via the pointwise strict limits
$\Phi(a)=\sum_{n=1}^{\infty}t_{n}\varphi^{(n)}(a)t_{n}^{*},\quad\mathbb{U}_{g}=\sum_{n=1}^{\infty}t_{n}\mathbbmss{u}^{(n)}_{g}t_{n}^{*}.$
Up to equivalence with a unitary in $\mathcal{M}(B)^{\beta}$, this cocycle representation does not depend on the choice of $(t_{n})_{n}$. In particular, in the special case that $(\varphi^{(n)},\mathbbmss{u}^{(n)})=(\varphi,\mathbbmss{u})$ for all $n$, we denote the resulting countable sum by $(\varphi^{\infty},\mathbbmss{u}^{\infty})$ or $(\varphi,\mathbbmss{u})^{\infty}$ and call it the infinite repeat of $(\varphi,\mathbbmss{u})$.
The last sentence of the following result and its justification follows the same line of reasoning as [8, Lemma 2.3] and [45, Lemma 6.1].
Proposition 1.7.
Let $(\varphi,\mathbbmss{u}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ be a cocycle representation. Then $(\varphi,\mathbbmss{u})$ is the infinite repeat of some cocycle representation if and only if there exists a sequence of isometries $r_{n}\in(\mathcal{M}(B)\cap\varphi(A)^{\prime}\cap\{\mathbbmss{u}_{g}\}_{g\in G}^{\prime})^{\beta}$ with $\sum_{n=1}^{\infty}r_{n}r_{n}^{*}=\mathbf{1}$ in the strict topology. If this is the case, then the $\mathrm{C}^{*}$-algebra $(\mathcal{M}(B)\cap\varphi(A)^{\prime})^{\beta^{\mathbbmss{u}}}$ has vanishing $K$-theory.
Proof.
The “if” part is trivial: If $r_{n}$ is such a sequence of isometries, then forming the infinite repeat of $(\varphi,\mathbbmss{u})$ via this chosen sequence yields the (on the nose) equality $(\varphi,\mathbbmss{u})=(\varphi,\mathbbmss{u})^{\infty}$.
For the “only if” part, suppose that $(\varphi,\mathbbmss{u})$ is the infinite repeat of some cocycle representation $(\psi,\mathbbmss{v})$. Let $t_{n}\in\mathcal{M}(B)^{\beta}$ be a sequence of isometries with $\sum_{n=1}^{\infty}t_{n}t_{n}^{*}=\mathbf{1}$ and such that
$\varphi(a)=\sum_{n=1}^{\infty}t_{n}\psi(a)t_{n}^{*},\quad a\in A,$
and
$\mathbbmss{u}_{g}=\sum_{n=1}^{\infty}t_{n}\mathbbmss{v}_{g}t_{n}^{*},\quad g\in G.$
Choose some sequence of pairwise disjoint infinite subsets $X_{j}\subset\mathbb{N}$ with $\mathbb{N}=\bigsqcup_{j=1}^{\infty}X_{j}$. For each $j\geq 1$ we choose a bijection $\kappa_{j}:\mathbb{N}\to X_{j}$ and consider the strict limit $r_{j}=\sum_{n=1}^{\infty}t_{\kappa_{j}(n)}t_{n}^{*}$. We see all these isometries commute with the range of both $\varphi$ and $\mathbbmss{u}$ pointwise and thus belong to the $\mathrm{C}^{*}$-algebra $(\mathcal{M}(B)\cap\varphi(A)^{\prime}\cap\{\mathbbmss{u}_{g}\}_{g\in G}^{\prime})^{\beta}$. Furthermore the sequence satisfies $\sum_{n=1}^{\infty}r_{n}r_{n}^{*}=\mathbf{1}$ in the strict topology.
For the last part of the statement, we keep a sequence $r_{n}$ chosen as in the statement. Consider the strict limit $r_{\infty}=\sum_{n=1}^{\infty}r_{n+1}r_{n}^{*}$, which is an isometry in $(\mathcal{M}(B)\cap\varphi(A)^{\prime})^{\beta^{\mathbbmss{u}}}$ and fits into the equation $r_{1}r_{1}^{*}+r_{\infty}r_{\infty}^{*}=\mathbf{1}$. For any given element $x\in(\mathcal{M}(B)\cap\varphi(A)^{\prime})^{\beta^{\mathbbmss{u}}}$, we can form its infinite repeat $x^{\infty}$ within $(\mathcal{M}(B)\cap\varphi(A)^{\prime})^{\beta^{\mathbbmss{u}}}$ via the sequence $(r_{n})_{n}$ and observe the equation
$x^{\infty}=\sum_{n=1}^{\infty}r_{n}xr_{n}^{*}=r_{1}xr_{1}^{*}+r_{\infty}\Big(\sum_{n=1}^{\infty}r_{n}xr_{n}^{*}\Big)r_{\infty}^{*}=x\oplus_{r_{1},r_{\infty}}x^{\infty}.$
Due to the additive properties of $K$-theory under Cuntz addition, this implies that the identity map of $(\mathcal{M}(B)\cap\varphi(A)^{\prime})^{\beta^{\mathbbmss{u}}}$ induces the zero morphism on its $K$-groups. This means that the $K$-groups vanish. ∎
Definition 1.8.
Let $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ be two cocycle representations. We write $(\varphi,\mathbbmss{u})\sim_{\mathrm{asymp}}(\psi,\mathbbmss{v})$, if there exists a norm-continuous path $U:[0,\infty)\to\mathcal{U}(\mathcal{M}(B))$ such that
  1. $\bullet$
    $\psi(a)=\displaystyle\lim_{t\to\infty}U_{t}\varphi(a)U_{t}^{*}$ for all $a\in A$;
  2. $\bullet$
    $\psi(a)-U_{t}\varphi(a)U_{t}^{*}\in B$ for all $a\in A$ and $t\geq 0$;
  3. $\bullet$
    $\lim_{t\to\infty}\max_{g\in K}\ \|\mathbbmss{v}_{g}-U_{t}\mathbbmss{u}_{g}\beta_{g}(U_{t})^{*}\|=0$ for all compact sets $K\subseteq G$;
  4. $\bullet$
    $\mathbbmss{v}_{g}-U_{t}\mathbbmss{u}_{g}\beta_{g}(U_{t})^{*}\in B$ for all $t\geq 0$ and $g\in G$.
If it is possible to choose $U$ to have its range in $\mathcal{U}(\mathbf{1}+B)$, then $(\varphi,\mathbbmss{u})$ and $(\psi,\mathbbmss{v})$ are called properly asymptotically unitarily equivalent. If it is additionally possible to arrange for $U_{0}=\mathbf{1}$, then $(\varphi,\mathbbmss{u})$ and $(\psi,\mathbbmss{v})$ are called strongly asymptotically unitarily equivalent.
Definition 1.9.
Let $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ be two cocycle representations.
  1. (i)
    We say that $(\varphi,\mathbbmss{u})$ absorbs $(\psi,\mathbbmss{v})$ if $(\varphi\oplus\psi,\mathbbmss{u}\oplus\mathbbmss{v})\sim_{\mathrm{asymp}}(\varphi,\mathbbmss{u})$.
  2. (ii)
    We say that $(\varphi,\mathbbmss{u})$ weakly absorbs $(\psi,\mathbbmss{v})$ if there exists a unitary $W\in\mathcal{U}(\mathcal{M}(B))$ such that $\varphi(a)-W(\varphi\oplus\psi)(a)W^{*}\in B$ for all $a\in A$ and $\mathbbmss{u}_{g}-W(\mathbbmss{u}_{g}\oplus\mathbbmss{v}_{g})\beta_{g}(W)^{*}\in B$ for all $g\in G$.
In line with the above condition, a given cocycle representation $(\varphi,\mathbbmss{u})$ is called absorbing if it absorbs every cocycle representation $(A,\alpha)\to(\mathcal{M}(B),\beta)$. It was proved as part of [21, Theorem 4.16] that such representations always exist as long as both $A$ and $B$ are separable.
Definition 1.10 ([21, Definition 4.3]).
Let $(\psi,\mathbbmss{v}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ be a cocycle representation, and let $\mathfrak{C}$ be a family of cocycle representations from $(A,\alpha)$ to $(\mathcal{M}(B),\beta)$. We say that $(\psi,\mathbbmss{v})$ is weakly contained in $\mathfrak{C}$, written $(\psi,\mathbbmss{v})\preccurlyeq\mathfrak{C}$, if the following is true.
For all compact sets $K\subseteq G$, $\mathcal{F}\subset A$, every $\varepsilon>0$, and every contraction $b\in B$ there exist $(\varphi^{(1)},\mathbbmss{u}^{(1)}),\dots,(\varphi^{(\ell)},\mathbbmss{u}^{(\ell)})\in\mathfrak{C}$ and a collection of elements
$\{c_{j,k}\mid j=1,\dots,N,\ k=1,\dots,\ell\}\subset B$
satisfying
$\max_{g\in K}\Big\|b^{*}\mathbbmss{v}_{g}\beta_{g}(b)-\sum_{k=1}^{\ell}\sum_{j=1}^{N}c_{j,k}^{*}\mathbbmss{u}^{(k)}_{g}\beta_{g}(c_{j,k})\Big\|\leq\varepsilon$
and
$\max_{a\in\mathcal{F}}\Big\|b^{*}\psi(a)b-\sum_{k=1}^{\ell}\sum_{j=1}^{N}c_{j,k}^{*}\varphi^{(k)}(a)c_{j,k}\Big\|\leq\varepsilon.$
In particular, if $\mathfrak{C}$ consists of a single cocycle representation $(\varphi,\mathbbmss{u})$, then we write $(\psi,\mathbbmss{v})\preccurlyeq(\varphi,\mathbbmss{u})$ in place of $(\psi,\mathbbmss{v})\preccurlyeq\left\{(\varphi,\mathbbmss{u})\right\}$, and in this case one can choose $\ell=1$ above.
Definition 1.11 ([21, Definition 4.5]).
Let $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ be two cocycle representations. We say that $(\psi,\mathbbmss{v})$ is contained in $(\varphi,\mathbbmss{u})$ at infinity, if the following is true. For all compact sets $K\subseteq G$, $\mathcal{F}\subset A$, every $\varepsilon>0$, and every contraction $b\in B$ there exists an element $x\in B$ satisfying
$\max_{g\in K}\|b^{*}\mathbbmss{v}_{g}\beta_{g}(b)-x^{*}\mathbbmss{u}_{g}\beta_{g}(x)\|\leq\varepsilon,\ \max_{a\in\mathcal{F}}\|b^{*}\psi(a)b-x^{*}\varphi(a)x\|\leq\varepsilon,\ \|x^{*}b\|\leq\varepsilon.$
Proposition 1.12 ([21, Lemma 4.11]).
Let $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ be two cocycle representations. The following are equivalent:
  1. (i)
    $(\psi,\mathbbmss{v})$ is contained in $(\varphi,\mathbbmss{u})$ at infinity;
  2. (ii)
    $(\varphi,\mathbbmss{u})$ absorbs $(\psi^{\infty},\mathbbmss{v}^{\infty})$.
  3. (iii)
    $(\varphi,\mathbbmss{u})$ weakly absorbs $(\psi^{\infty},\mathbbmss{v}^{\infty})$.
In particular, $(\varphi,\mathbbmss{u})$ contains itself at infinity if and only if modulo $B$, it can be expressed as the unitary conjugate of an infinite repeat.
1.2. The Cuntz–Thomsen picture of equivariant $KK$-theory
Definition 1.13 (cf. [42, Section 3]).
Let $\alpha:G\curvearrowright A$ and $\beta:G\curvearrowright B$ be two actions on $\mathrm{C}^{*}$-algebras such that $A$ is separable and $B$ is $\sigma$-unital. An (equivariant) $(\alpha,\beta)$-Cuntz pair is a pair of cocycle representations
$(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v}):(A,\alpha)\to(\mathcal{M}(B\otimes\mathcal{K}),\beta\otimes\operatorname{id}_{\mathcal{K}}),$
such that the pointwise differences $\varphi-\psi$ and $\mathbbmss{u}-\mathbbmss{v}$ take values in $B\otimes\mathcal{K}$. [^4] In Thomsen’s article it was also assumed that the map $\mathbbmss{u}-\mathbbmss{v}$ is norm-continuous. This turns out to be redundant, see [41, Proposition 6.9]. A pair of the specific form $\big((\varphi,\mathbbmss{u}),(\varphi,\mathbbmss{u})\big)$ is called degenerate. Whenever $\beta$ is assumed to be strongly stable (this is our default assumption), we also allow $(B,\beta)$ in place of $(B\otimes\mathcal{K},\beta\otimes\operatorname{id}_{\mathcal{K}})$ appearing in the definition of an equivariant $(\alpha,\beta)$-Cuntz pair.
Notation 1.14.
Given a $\mathrm{C}^{*}$-algebra $B$, we denote $B[0,1]=\mathcal{C}[0,1]\otimes B$. If one has an action $\beta:G\curvearrowright B$, we consider the induced $G$-action on $B[0,1]$ given by $\beta[0,1]=\operatorname{id}_{\mathcal{C}[0,1]}\otimes\beta$.
Definition 1.15 (see [42, Section 3] and [21, Section 2]).
Let $A$ be a separable $\mathrm{C}^{*}$-algebra and $B$ a $\sigma$-unital $\mathrm{C}^{*}$-algebra. For two actions $\alpha:G\curvearrowright A$ and $\beta:G\curvearrowright B$, let $\mathbb{E}^{G}(\alpha,\beta)$ denote the set of all $(\alpha,\beta)$-Cuntz pairs.
Two elements $\big((\varphi^{0},\mathbbmss{u}^{0}),(\psi^{0},\mathbbmss{v}^{0})\big)$ and $\big((\varphi^{1},\mathbbmss{u}^{1}),(\psi^{1},\mathbbmss{v}^{1})\big)$ in $\mathbb{E}^{G}(\alpha,\beta)$ are called homotopic, abbreviated $\big((\varphi^{0},\mathbbmss{u}^{0}),(\psi^{0},\mathbbmss{v}^{0})\big)\sim_{h}\big((\varphi^{1},\mathbbmss{u}^{1}),(\psi^{1},\mathbbmss{v}^{1})\big)$, if there exists an $(\alpha,\beta[0,1])$-Cuntz pair that restricts to $\big((\varphi^{0},\mathbbmss{u}^{0}),(\psi^{0},\mathbbmss{v}^{0})\big)$ upon evaluation at $0\in[0,1]$, and restricts to $\big((\varphi^{1},\mathbbmss{u}^{1}),(\psi^{1},\mathbbmss{v}^{1})\big)$ upon evaluation at $1\in[0,1]$. An $(\alpha,\beta)$-Cuntz pair of the form $\big((\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v})\big)$ with $\varphi=\psi=0$ is called a cocycle pair and is denoted by $(\mathbbmss{u},\mathbbmss{v})$ with slight abuse of notation. We define $\mathbb{E}_{0}^{G}(\alpha,\beta)$ as the set of all anchored $(\alpha,\beta)$-Cuntz pairs, i.e., those $\big((\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v})\big)\in\mathbb{E}^{G}(\alpha,\beta)$ such that $(\mathbbmss{u},\mathbbmss{v})\sim_{h}(\mathbf{1},\mathbf{1})$.
For any unital inclusion $\mathcal{O}_{2}\subseteq\mathcal{M}(B\otimes\mathcal{K})^{\beta\otimes\operatorname{id}_{\mathcal{K}}}$ with generating isometries $t_{1},t_{2}$, one can perform the Cuntz addition for two $(\alpha,\beta)$-Cuntz pairs as
[…]
This Cuntz pair is independent of the choice of $t_{1},t_{2}$ up to homotopy; see [42, Lemma 3.4].
Remark 1.16 (see [21, Proposition 2.12]).
The quotient $\mathbb{E}^{G}(\alpha,\beta)/{\sim_{h}}$ becomes an abelian group with Cuntz addition. The homotopy classes of cocycle pairs form a canonical subgroup $H_{\beta}$ independent of $\alpha$. It was proved by Thomsen in [42, Theorem 3.5] that the group quotient of $\mathbb{E}^{G}(\alpha,\beta)/{\sim_{h}}$ modulo $H_{\beta}$ is naturally isomorphic to $KK^{G}(\alpha,\beta)$ if one defines the latter via Kasparov’s original approach [31]. For an $(\alpha,\beta)$-Cuntz pair consisting of $(\varphi,\mathbbmss{u})$ and $(\psi,\mathbbmss{v})$, we denote its associated homotopy class by $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v})]$. Under this identification, one has that the inclusion map $\mathbb{E}^{G}_{0}(\alpha,\beta)\subseteq\mathbb{E}^{G}(\alpha,\beta)$ also induces a natural isomorphism of abelian groups $\mathbb{E}^{G}_{0}(\alpha,\beta)/{\sim_{h}}\cong KK^{G}(\alpha,\beta)$. In other words, $KK^{G}(\alpha,\beta)$ may be defined as the abelian group of homotopy classes of anchored $(\alpha,\beta)$-Cuntz pairs.
2. $K_{1}$-injectivity of certain auxiliary $\mathrm{C}^{*}$-algebras
The following notation was already used in [21] and is inspired by notation appearing in the work of Dadarlat–Eilers in [12, Section 3]. We note that for $\mathrm{C}^{*}$-dynamics, a very similar type of auxiliary $\mathrm{C}^{*}$-algebras were considered by Thomsen in [45, Section 6] and their conceptual role was similar to how we will use the objects below in this article.
Notation 2.1.
Let $(\varphi,\mathbbmss{u}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ be a cocycle representation. We then consider the two $\mathrm{C}^{*}$-algebras
$\mathscr{D}_{(\varphi,\mathbbmss{u})}=\left\{x\in\mathcal{M}(B)\mid[x,\varphi(A)]\subseteq B,\left\{x-\beta^{\mathbbmss{u}}_{g}(x)\right\}_{g\in G}\subseteq B\right\}$
and
$\mathscr{D}^{w}_{(\varphi,\mathbbmss{u})}=\left\{x\in\mathcal{M}(B)\mid[x,\varphi(A)]\subseteq B,\left\{\varphi(a)(\beta^{\mathbbmss{u}}_{g}(x)-x)\right\}_{a\in A,g\in G}\subseteq B\right\}.$
If $\varphi:A\to\mathcal{M}(B)$ is any *-homomorphism, then it can be viewed as a cocycle representation for actions of the trivial group, and we write
$\mathscr{D}_{\varphi}=\left\{x\in\mathcal{M}(B)\mid[x,\varphi(A)]\subseteq B\right\}.$
Remark 2.2.
Let $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ be two cocycle representations. Suppose there exists $U\in\mathcal{U}(\mathcal{M}(B))$ such that $(\varphi,\mathbbmss{u})$ agrees with $\operatorname{Ad}(U)\circ(\psi,\mathbbmss{v})$ modulo $B$. Based on the definitions, we can directly conclude that $\operatorname{Ad}(U)\in\operatorname{Aut}(\mathcal{M}(B))$ restricts to isomorphisms
$\mathscr{D}^{w}_{(\psi,\mathbbmss{v})}\cong\mathscr{D}^{w}_{(\varphi,\mathbbmss{u})}\quad\text{and}\quad\mathscr{D}_{(\psi,\mathbbmss{v})}\cong\mathscr{D}_{(\varphi,\mathbbmss{u})}.$
If we assume that $(\varphi,\mathbbmss{u})$ contains itself at infinity and apply subsection 1.1, it follows that one has isomorphisms $\mathscr{D}^{w}_{(\varphi,\mathbbmss{u})}\cong\mathscr{D}^{w}_{(\varphi,\mathbbmss{u})^{\infty}}$ and $\mathscr{D}_{(\varphi,\mathbbmss{u})}\cong\mathscr{D}_{(\varphi,\mathbbmss{u})^{\infty}}$.
For what follows, we recall that a unital $\mathrm{C}^{*}$-algebra $D$ is called properly infinite if it contains two isometries $s,t\in D$ with $s^{*}t=0$.
Notation 2.3 (see [32, Subsection 4.2]).
Let $D$ be a properly infinite unital $\mathrm{C}^{*}$-algebra. We say that a projection $p\in D$ is a splitting projection, if there exist isometries $s,t\in D$ such that $ss^{*}\leq p$ and $tt^{*}\leq\mathbf{1}-p$. We say that $D$ is in Cuntz standard form, if there exists a unital inclusion $\mathcal{O}_{2}\subseteq D$.
Remark 2.4.
It follows from Cuntz’s fundamental article [7] that the canonical map $p\mapsto[p]_{0}$ induces a one-to-one correspondence between unitary equivalence classes of splitting projections in $D$ and the group $K_{0}(D)$. Furthermore, if $D$ is in Cuntz standard form and $p\in D$ is a splitting projection, then $p]_{0}=0$ if and only if $p\sim_{\mathrm{MvN}}\mathbf{1}-p\sim_{\mathrm{MvN}}\sim\mathbf{1}$. Note that in [11], projections with the latter property were referred to as “proper” projections. We shall not use this terminology because it may be considered too ambiguous from today’s point of view, but wish to point this out to the reader as we will reference said article below.
Lemma 2.5.
Let $D$ be a properly infinite unital $\mathrm{C}^{*}$-algebra in Cuntz standard form. Let $p,q\in D$ be two splitting projections with vanishing $K_{0}$-class. Suppose that there exist isometries $s_{1},s_{2}\in D$ such that $s_{1}s_{1}^{*}\leq p$, $s_{2}s_{2}^{*}\leq q$ and $\|s_{1}^{*}s_{2}\|<1$. Then $p$ and $q$ are homotopic.
Proof.
As both $p$ and $q$ are splitting projections, it immediately follows for $j=1,2$ that the range projections $s_{j}s_{j}^{*}$ are also splitting projections with vanishing $K_{0}$-class. Using Cuntz–Higson’s [11, Lemma 1] repeatedly, it follows that we have a chain of homotopies
$p\sim_{h}\mathbf{1}-p\sim_{h}s_{1}s_{1}^{*}\sim_{h}s_{2}s_{2}^{*}\sim_{h}\mathbf{1}-q\sim_{h}q.$
This finishes the argument. ∎
Recall that a unital $\mathrm{C}^{*}$-algebra $A$ is called $K_{1}$-injective if the canonical map $\mathcal{U}(A)/\mathcal{U}_{0}(A)\to K_{1}(A)$ is injective. The following serves as our main criterion for $K_{1}$-injectivity of properly infinite $\mathrm{C}^{*}$-algebras.
Lemma 2.6.
Let $D$ be a properly infinite unital $\mathrm{C}^{*}$-algebra. The following are equivalent:
  1. (i)
    $D$ is $K_{1}$-injective.
  2. (ii)
    For every pair of splitting projections $p,q\in D$, one has $p\sim_{\mathrm{MvN}}q$ if and only if $p\sim_{h}q$.
  3. (iii)
    There exists some splitting projection $p\in D$ such that for every splitting projection $q\in D$, one has $p\sim_{\mathrm{MvN}}q$ if and only if $p\sim_{h}q$.
Proof.
[…]
The statement below is the consequence of a standard functional calculus argument. We shall omit its easy proof.
Proposition 2.7.
For every $\varepsilon>0$, there exists a constant $\delta>0$ with the following property. Let $A$ be any $\mathrm{C}^{*}$-algebra and $a\in A$ a positive contraction.
  1. $\bullet$
    If $x\in A$ is any contraction with $\|[x,a]\|<\delta$, then $\|[x,\sqrt{a}]\|<\varepsilon$.
  2. $\bullet$
    If $\alpha\in\operatorname{Aut}(A)$ is any automorphism with $\|a-\alpha(a)\|<\delta$, then $\|\sqrt{a}-\alpha(\sqrt{a})\|<\varepsilon$.
The following is the main technical result of this section. A special case of it, stated after it, is the main ingredient to answer Introduction affirmatively, and its general form enables many of the results in the rest of the article.
Theorem 2.8.
Let $\alpha:G\curvearrowright A$ be an action on a separable $\mathrm{C}^{*}$-algebra and $\beta:G\curvearrowright B$ a strongly stable action on a $\sigma$-unital $\mathrm{C}^{*}$-algebra. Let $(\varphi,\mathbbmss{u}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ be a cocycle representation that contains itself at infinity. Let $\mathcal{E}$ be a unital $\mathrm{C}^{*}$-algebra for which one of the following holds:
  1. (1)
    There exists an ideal $\mathcal{J}\triangleleft\mathscr{D}_{(\varphi,\mathbbmss{u})}^{w}$ such that $\mathcal{E}\cong\mathscr{D}_{(\varphi,\mathbbmss{u})}^{w}/\mathcal{J}$.
  2. (2)
    There exists an ideal $\mathcal{J}\triangleleft\mathscr{D}_{(\varphi,\mathbbmss{u})}$ such that $\mathcal{E}\cong\mathscr{D}_{(\varphi,\mathbbmss{u})}/\mathcal{J}$.
Then $\mathcal{E}$ is $K_{1}$-injective.
Proof.
By section 2, we may assume without loss of generality that $(\varphi,\mathbbmss{u})$ is an infinite repeat. By subsection 1.1, we may assume that there exists a sequence of isometries $r_{n}\in(\mathcal{M}(B)\cap\varphi(A)^{\prime})^{\beta^{\mathbbmss{u}}}$ with $\mathbf{1}=\sum_{n=1}^{\infty}r_{n}r_{n}^{*}$. Depending on which case is proved, we write either $\mathscr{D}=\mathscr{D}_{(\varphi,\mathbbmss{u})}$ or $\mathscr{D}=\mathscr{D}_{(\varphi,\mathbbmss{u})}^{w}$, so that $\mathcal{E}=\mathscr{D}/\mathcal{J}$ in both cases. We shall assume that $\mathcal{E}$ is non-zero, otherwise there is nothing to prove.
We consider $p=P+\mathcal{J}$ for $P=\sum_{n=1}^{\infty}r_{2n}r_{2n}^{*}\in\mathscr{D}$, which defines a splitting projection in $\mathcal{E}$ with vanishing $K_{0}$-class (as $P$ is already such an element in $\mathscr{D}$). Let $q\in\mathcal{E}$ be another splitting projection with vanishing $K_{0}$-class. By section 2, the claim amounts to showing that $p$ and $q$ are homotopic. By section 2, $q$ is uniquely determined modulo unitary equivalence, thus there exists a unitary $w\in\mathcal{E}$ such that $w(\mathbf{1}-p)w^{*}=q$. Let $W\in\mathscr{D}$ be a contraction with $w=W+\mathcal{J}$. Set $Q=W(\mathbf{1}-P)W^{*}=W\big(\sum_{n=1}^{\infty}r_{2n-1}r_{2n-1}^{*}\big)W^{*}$, which then satisfies $q=Q+\mathcal{J}$.
We apply section 1 and choose an increasing approximate unit $(e_{n})_{n\geq 1}$ in $B$ with $e_{n+1}e_{n}=e_{n}$ that is $\varphi(A)$-quasicentral and approximately $\beta^{\mathbbmss{u}}$-invariant. Let $K_{n}\subseteq G$ be an increasing sequence of compact sets with union $G$ and $F_{n}\subseteq A$ an increasing sequence of finite subsets of contractions whose union is dense in the unit ball of $A$. For every $n\geq 1$, we apply section 2 to find a universal constant $\delta_{n}>0$ that satisfies the conclusion therein for $2^{-(n+2)}$ in place of $\varepsilon$. We assume $\delta_{n}\leq 2^{-n}$ for each $n\geq 1$ and that the sequence $(\delta_{n})_{n}$ is decreasing. We begin by choosing an even number $m_{1}\geq 1$ with
$\max_{g\in K_{1}}\|e_{n}-\beta_{g}^{\mathbbmss{u}}(e_{n})\|+\max_{a\in F_{1}}\|[e_{n},\varphi(a)]\|<\delta_{1}/2\ ,\quad n\geq m_{1}.$
Set $f_{1}=r_{m_{1}}e_{m_{1}}r_{m_{1}}^{*}$. As $r_{n}r_{n}^{*}\to 0$ strictly as $n\to\infty$, we may find an odd number $n_{1}>m_{1}$ such that $\|e_{m_{1}}r_{m_{1}}^{*}Wr_{n_{1}}\|<\frac{1}{4}$ and set $h_{1}=Wr_{n_{1}}e_{n_{1}}r_{n_{1}}^{*}W^{*}$. Next, choose an even number $m_{2}>n_{1}$ with
$\max_{g\in K_{2}}\|e_{n}-\beta_{g}^{\mathbbmss{u}}(e_{n})\|+\max_{a\in F_{2}}\|[e_{n},\varphi(a)]\|<\delta_{2}/2,\quad n\geq m_{2},$
and
$\|r_{n}^{*}Wr_{n_{1}}e_{n_{1}}\|<\frac{1}{8},\quad n\geq m_{2}.$
Set $f_{2}=r_{m_{2}}e_{m_{2}}r_{m_{2}}^{*}$. Then find an odd number $n_{2}>m_{2}$ such that
$\|e_{m_{1}}r_{m_{1}}^{*}Wr_{n_{2}}\|<\frac{1}{8},\quad\|e_{m_{2}}r_{m_{2}}^{*}Wr_{n_{2}}\|<\frac{1}{16},$
and set $h_{2}=Wr_{n_{2}}e_{n_{2}}r_{n_{2}}^{*}W^{*}$.
By repeating these steps inductively, we end up with increasing sequences of (alternatingly even and odd) natural numbers $m_{1}<n_{1}<m_{2}<n_{2}<\dots$ such that
(e2.1) $\sup_{n\geq m_{k}}\Big(\max_{g\in K_{k}}\|e_{n}-\beta_{g}^{\mathbbmss{u}}(e_{n})\|+\max_{a\in F_{k}}\|[e_{n},\varphi(a)]\|\Big)<\delta_{k}/2,\quad k\geq 1,$
and the positive contractions
$f_{k}=r_{m_{k}}e_{m_{k}}r_{m_{k}}^{*},\quad h_{k}=Wr_{n_{k}}e_{n_{k}}r_{n_{k}}^{*}W^{*}$
satisfy the inequality $\|f_{k}h_{j}\|<2^{-(j+k)}$ for all $j,k\geq 1$.
Consider the strict limits $\mathfrak{F}=\sum_{k=1}^{\infty}f_{k}$ and $\mathfrak{H}=\sum_{k=1}^{\infty}h_{k}$. Clearly $\mathfrak{F}\leq P$ and $\mathfrak{H}\leq Q$ and the above inequality implies $\|\mathfrak{F}\cdot\mathfrak{H}\|\leq\sum_{j,k=1}^{\infty}\|f_{k}h_{j}\|<1$.
For notational convenience, set $e_{0}:=0=:e_{-1}$ and $m_{0}=n_{0}=0=m_{-1}=n_{-1}$. For $k\geq 1$, define $a_{k}=(e_{m_{k-1}}-e_{m_{k-2}})^{1/2}$ and $b_{k}=(e_{n_{k-1}}-e_{n_{k-2}})^{1/2}$. As $(e_{n})_{n}$ was an approximate unit, note that
$\sum_{k=1}^{\infty}a_{k}^{2}=\lim_{\ell\to\infty}e_{m_{\ell-1}}=\mathbf{1}=\lim_{\ell\to\infty}e_{n_{\ell-1}}=\sum_{k=1}^{\infty}b_{k}^{2}.$
By condition (e2.1), we have for all $k\geq 3$ that
$\max_{g\in K_{k-2}}\|(e_{m_{k-1}}-e_{m_{k-2}})-\beta_{g}^{\mathbbmss{u}}(e_{m_{k-1}}-e_{m_{k-2}})\|\leq\delta_{k-2}$
and
$\max_{a\in F_{k-2}}\|[(e_{m_{k-1}}-e_{m_{k-2}}),\varphi(a)]\|\leq\delta_{k-2}.$
By our choice of the constants $\delta_{k}$, it follows for all $k\geq 3$ that
(e2.2) $\max_{g\in K_{k-2}}\|a_{k}-\beta_{g}^{\mathbbmss{u}}(a_{k})\|\leq 2^{-k},\quad\max_{a\in F_{k-2}}\|[a_{k},\varphi(a)]\|\leq 2^{-k}.$
Analogously, we get for all $k\geq 3$ that
(e2.3) $\max_{g\in K_{k-2}}\|b_{k}-\beta_{g}^{\mathbbmss{u}}(b_{k})\|\leq 2^{-k},\quad\max_{a\in F_{k-2}}\|[b_{k},\varphi(a)]\|\leq 2^{-k}.$
Consider the strict limit $R_{1}=\sum_{k=1}^{\infty}r_{m_{k}}a_{k}$. It is an isometry since
$R_{1}^{*}R_{1}=\sum_{k,j=1}^{\infty}a_{k}r_{m_{k}}^{*}r_{m_{j}}a_{j}=\sum_{k=1}^{\infty}a_{k}^{2}=\mathbf{1}.$
We also have
$\mathfrak{F}R_{1}=\sum_{k,j=1}^{\infty}r_{m_{j}}e_{m_{j}}r_{m_{j}}^{*}r_{m_{k}}a_{k}=\sum_{k=1}^{\infty}r_{m_{k}}\underbrace{e_{m_{k}}a_{k}}_{=a_{k}}=R_{1}.$
Condition (e2.2) implies for all $a\in\bigcup_{n}F_{n}$ that
$R_{1},\varphi(a)]=\sum_{k=1}^{\infty}r_{m_{k}}[a_{k},\varphi(a)]\in B$
and for all $g\in G$ that
$R_{1}-\beta^{\mathbbmss{u}}_{g}(R_{1})=\sum_{k=1}^{\infty}r_{m_{k}}\big(a_{k}-\beta^{\mathbbmss{u}}_{g}(a_{k})\big)\in B,$
and hence $R_{1}\in\mathscr{D}_{(\varphi,\mathbbmss{u})}\subseteq\mathscr{D}$.
Likewise we can consider the strict limit $R_{2}^{\prime}=\sum_{k=1}^{\infty}r_{n_{k}}b_{k}$ and $R_{2}=WR_{2}^{\prime}$. As above, condition (e2.3) implies $R_{2}^{\prime}\in\mathscr{D}_{(\varphi,\mathbbmss{u})}\subseteq\mathscr{D}$ and thus $R_{2}\in\mathscr{D}$. Exactly as above it follows that $R_{2}^{\prime}$ is an isometry and hence
$R_{2}^{*}R_{2}-\mathbf{1}={R_{2}^{\prime}}^{*}(W^{*}W-\mathbf{1})R_{2}^{\prime}\in\mathcal{J}.$
Here we used that $W$ was chosen as a lift of the unitary $w\in\mathcal{E}$. Furthermore,
$\begin{array}[]{cll}\mathfrak{H}R_{2}&=&W\Big(\sum_{k=1}^{\infty}r_{n_{k}}e_{n_{k}}r_{n_{k}}^{*}\Big)W^{*}WR_{2}^{\prime}\\ &\equiv_{\mathcal{J}}&W\Big(\sum_{k=1}^{\infty}r_{n_{k}}e_{n_{k}}r_{n_{k}}^{*}\Big)R_{2}^{\prime}\\ &=&W\sum_{k,j=1}^{\infty}r_{n_{j}}e_{n_{j}}r_{n_{j}}^{*}r_{n_{k}}b_{k}\\ &=&W\sum_{k=1}^{\infty}r_{n_{k}}\underbrace{e_{n_{k}}b_{k}}_{=b_{k}}\\ &=&WR_{2}^{\prime}\ =\ R_{2}.\end{array}$
To summarize, if we set $s_{j}=R_{j}+\mathcal{J}$ for $j=1,2$, then we have found two isometries $s_{1},s_{2}\in\mathcal{E}$ with
$s_{1}^{*}s_{1}^{*}\leq\mathfrak{F}+\mathcal{J}\leq P+\mathcal{J}=p,\quad s_{2}s_{2}^{*}\leq\mathfrak{H}+\mathcal{J}\leq Q+\mathcal{J}=q,$
and
$\|s_{1}^{*}s_{2}\|\leq\|R_{1}^{*}\mathfrak{F}\mathfrak{H}R_{2}\|\leq\|\mathfrak{F}\mathfrak{H}\|<1.$
With section 2, we may conclude that $p$ and $q$ are homotopic in $\mathcal{E}$. ∎
As a special case of this result, we obtain a positive solution to the $KK$-uniqueness problem for $\mathrm{C}^{*}$-algebras posed in recent work of Carrión et al; see also the thorough discussion following [39, Problem LXII]. Note that the special case $A=0$ recovers [19, Proposition 4.9], the special case $B=\mathcal{K}$ recovers [37, Lemma 3(2)], and [35, Theorem 2.5, 2.9], [36, Theorem 3.28] as well as [28, Corollary 4.2] are recovered as various other special cases. We note that the unitality condition written in brackets is an optional extra assumption one may add to the theorem in case $A$ is unital, so we obtain two versions of the statement below, with or without units. In particular we also obtain Introduction.
Corollary 2.9 (see [6, Question 5.17] and the surrounding discussion).
Let $A$ be a separable $\mathrm{C}^{*}$-algebra and $B$ a $\sigma$-unital stable $\mathrm{C}^{*}$-algebra. Let $\varphi:A\to\mathcal{M}(B)$ be a (unitally) absorbing representation and let $\bar{\varphi}:A\to\mathcal{Q}(B)=\mathcal{M}(B)/B$ be its induced homomorphism to the corona algebra. Then the relative commutant $\mathcal{Q}(B)\cap\bar{\varphi}(A)^{\prime}$ is $K_{1}$-injective. The same conclusion holds if $\varphi$ is instead assumed to be weakly nuclear and nuclearly (unitally) absorbing. In particular, Introduction has an affirmative answer.
Proof.
We apply section 2 to the case $G=\{1\}$ with trivial actions on $A$ and $B$. Since $\varphi$ is either a (unitally) absorbing representation or is weakly nuclear and nuclearly (unitally) absorbing, it absorbs its infinite repeat $\varphi^{\infty}$, so it contains itself at infinity. If one keeps in mind that $B$ sits in $\mathscr{D}_{\varphi}$ as an ideal and compares definitions, one has $\mathcal{Q}(B)\cap\bar{\varphi}(A)^{\prime}=\mathscr{D}_{\varphi}/B$. So if we apply the conclusion of section 2 with $\mathcal{J}=B$, the claim follows. ∎
In the subsequent sections, we will generalize this consequence related to $KK$-uniqueness to the setting of generalized equivariant $KK$-groups with an approach that can be considered self-contained if one combines the methodology of this article with that of [21].
3. On a generalized equivariant $KK$-group construction
Definition 3.1.
Let $\mathfrak{C}$ be a set of cocycle representations from $(A,\alpha)$ to $(\mathcal{M}(B),\beta)$. We say that $\mathfrak{C}$ is downward closed, if for every cocycle representation $(\varphi,\mathbbmss{u})$, we have that $(\varphi,\mathbbmss{u})\preccurlyeq\mathfrak{C}$ implies $(\varphi,\mathbbmss{u})\in\mathfrak{C}$. We say that $\mathfrak{C}$ is an eligible set of cocycle representations for $(\alpha,\beta)$, if it is downward closed and contains every cocycle representation of the form $(0,\mathbbmss{u})$, where $\mathbbmss{u}$ is an arbitrary $\beta$-cocycle in $\mathcal{M}(B)$.
Remark 3.2.
Note that in the non-equivariant situation, i.e. $G=\{1\}$, being eligible is the same as being downward closed. In general, if $\mathfrak{C}$ is a downward closed set as above, then it follows from [21, Proposition 4.7] that $\mathfrak{C}$ is $\sigma$-additive in the sense of ‘Definition 4.13’ therein, i.e., it is closed under countable sums as considered in subsection 1.1. If $B$ is separable, then it follows by [21, Theorem 4.16] that $\mathfrak{C}$ always has an absorbing element, i.e., an element $(\varphi,\mathbbmss{u})\in\mathfrak{C}$ that absorbs every element of $\mathfrak{C}$ in the sense of subsection 1.1.
Definition 3.3.
Let $\mathfrak{C}$ be a set of cocycle representations from $(A,\alpha)$ to $(\mathcal{M}(B),\beta)$. We define $\mathfrak{C}[0,1]$ as the set of cocycle representations $(\Phi,\mathbb{U}):(A,\alpha)\to(\mathcal{M}(B[0,1]),\beta[0,1])$ such that $\operatorname{ev}_{t}\circ(\Phi,\mathbb{U})=:(\Phi_{t},\mathbb{U}_{t})\in\mathfrak{C}$ for all $t\in[0,1]$, where $\operatorname{ev}_{t}:B[0,1]\to B$ is the evaluation map at $t$.
Definition 3.4.
Suppose that $\mathfrak{C}$ is an eligible set of cocycle representations from $(A,\alpha)$ to $(\mathcal{M}(B),\beta)$. An (equivariant) $(\mathfrak{C};\alpha,\beta)$-Cuntz pair is an equivariant $(\alpha,\beta)$-Cuntz pair consisting of cocycle representations $(\varphi,\mathbbmss{u})$ and $(\psi,\mathbbmss{v})$ belonging to $\mathfrak{C}$. We define $\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)$ to be the set of all such pairs and $\mathbb{E}_{0}^{G}(\mathfrak{C};\alpha,\beta)$ to be the set of all such pairs that are also anchored in the sense of subsection 1.2.
We say that two elements $x_{0},x_{1}\in\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)$ are $\mathfrak{C}$-homotopic, written $x_{0}\sim_{h,\mathfrak{C}}x_{1}$, if there exists $X\in\mathbb{E}^{G}(\mathfrak{C}[0,1];\alpha,\beta[0,1])$ with $\operatorname{ev}_{j}\circ X=x_{j}$ for j=0,1j=0,1.555For the sake of notational clarity we slightly abuse notation here. The Cuntz pair $\operatorname{ev}_{j}\circ X$ is meant to be the Cuntz pair coming from composing both components of $X$ from the left with the evaluation map in $j$. Given a pair $\big((\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v})\big)\in\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)$, we denote by $\big[(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v})\big]$ its $\mathfrak{C}$-homotopy class. As was the case before, the subset $\mathbb{E}^{G}_{0}(\mathfrak{C};\alpha,\beta)\subseteq\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)$ is closed under $\mathfrak{C}$-homotopy.
Remark 3.5.
Exactly as one does in the Cuntz–Thomsen picture of equivariant $KK$-theory, one observes in the context of section 3 that Cuntz addition of $(\mathfrak{C};\alpha,\beta)$-Cuntz pairs descends to a well-defined abelian group structure on $\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)/_{\sim_{h,\mathfrak{C}}}$. The proof of this fact (see for instance [21, Lemma 2.10], though the argument goes back much further than that) carries over word for word from the (known) case of $\mathfrak{C}$ being the set of all cocycle representations. Here the fact that $\mathfrak{C}$ is downward closed ensures that $\mathfrak{C}$ is closed under Cuntz addition and unitary equivalence, which enables the known proof to work here. The fact that $\mathfrak{C}$ contains all cocycle representations of the form $(0,\mathbbmss{u})$ for arbitrary $\beta$-cocycles $\mathbbmss{u}$ implies that there is a canonical homomorphism from the group $H_{\beta}$ (see subsection 1.2) to $\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)/_{\sim_{h,\mathfrak{C}}}$. On the other hand, the assignment $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v})]\mapsto[(0,\mathbbmss{u}),(0,\mathbbmss{v})]$ yields a well-defined homomorphism $\mathfrak{f}:\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)/_{\sim_{h,\mathfrak{C}}}\to H_{\beta}$ that is a left inverse of the inclusion map $H_{\beta}\to\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)/_{\sim_{h,\mathfrak{C}}}$. We call $\mathfrak{f}$ the forgetful map and observe that $\ker(\mathfrak{f})=\mathbb{E}^{G}_{0}(\mathfrak{C};\alpha,\beta)/_{\sim_{h,\mathfrak{C}}}$. This is analogous to the observations made in [21, Proposition 2.12].
Definition 3.6.
Suppose that $\mathfrak{C}$ is an eligible set of cocycle representations from $(A,\alpha)$ to $(\mathcal{M}(B),\beta)$. We define $KK^{G}(\mathfrak{C};\alpha,\beta)$ to be the quotient of abelian groups $\big(\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)/_{\sim_{h,\mathfrak{C}}}\big)/H_{\beta}$. By section 3, the forgetful map induces a direct sum decomposition $\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)/_{\sim_{h,\mathfrak{C}}}\cong H_{\beta}\oplus\mathbb{E}_{0}^{G}(\mathfrak{C};\alpha,\beta)/_{\sim_{h,\mathfrak{C}}}$ and therefore one has a natural isomorphism $KK^{G}(\mathfrak{C};\alpha,\beta)\cong\mathbb{E}_{0}^{G}(\mathfrak{C};\alpha,\beta)/_{\sim_{h,\mathfrak{C}}}$.
Example 3.7.
Let us have a brief discussion in what way section 3 extends various known constructions.
  1. (1)
    If $\mathfrak{C}$ denotes the set of all cocycle representations, then $KK^{G}(\mathfrak{C};\alpha,\beta)$ agrees with the equivariant $KK$-group $KK^{G}(\alpha,\beta)$ on the nose in Thomsen’s description [42].
  2. (2)
    For $G=\{1\}$, we may choose as $\mathfrak{C}$ the set of all weakly nuclear representations $A\to\mathcal{M}(B)$. Then $KK(\mathfrak{C};A,B)$ is naturally isomorphic to Skandalis’ nuclear KKKK-group $KK_{\mathrm{nuc}}(A,B)$ [40]; see [18, Section 12] for a rigorous treatment of the Cuntz picture in that setting.
  3. (3)
    Still assuming $G=\{1\}$, Kirchberg’s ideal-related $KK$-groups fall under the above framework. We refer to [18, Sections 10–12] for a rigorous treatment of the concepts mentioned here. For a topological space $X$, recall that an action of $X$ on a $\mathrm{C}^{*}$-algebra $A$ consists of an order preserving map $\Phi_{A}:\mathcal{O}(X)\to\mathcal{I}(A)$, where the domain is the lattice of open subsets of $X$ and the codomain is the ideal lattice of $A$. For an open set $U\subseteq X$, one writes $A(U)=\Phi_{A}(U)$ as short-hand. A $\mathrm{C}^{*}$-algebra with a given action of $X$ is called an $X$-$\mathrm{C}^{*}$-algebra. If we are given $X$-$\mathrm{C}^{*}$-algebras $(A,\Phi_{A})$ and $(B,\Phi_{B})$, then a *-homomorphism $\varphi:A\to\mathcal{M}(B)$ is called weakly $X$-equivariant if for all $b\in B$, $a\in A$ and every open set $U\subseteq X$ one has $b^{*}\varphi(A(U))b\subseteq B(U)$. It is immediate that if $A$ is separable and $B$ is stable and $\sigma$-unital, then the set $\mathfrak{C}_{X}$ of all weakly $X$-equivariant representations $A\to\mathcal{M}(B)$ is an eligible set. Upon comparing with the literature, we see that the group $KK(\mathfrak{C}_{X};A,B)$ in the sense of section 3 agrees with Kirchberg’s group $KK(X;A,B)$; see [18, Proposition 2.28]. The same is true if we consider the subset of representations $\mathfrak{C}_{X}^{\mathrm{nuc}}\subseteq\mathfrak{C}_{X}$ that are additionally weakly nuclear, in which case we obtain the nuclear ideal-related group $KK_{\mathrm{nuc}}(X;A,B)$.
  4. (4)
    Gabe’s article [18] includes a more detailed discussion about the $KK$-theory of $\mathcal{C}_{0}(X)$-algebras when $X$ is locally compact Hausdorff, which can be interpreted as the $KK$-theory of $X$-$\mathrm{C}^{*}$-algebras. From this point of view, if now $G\neq\{1\}$ and $X$ is compact Hausdorff, then Kasparov’s notion [31] of $KK$-theory for $G$-$\mathcal{C}(X)$-algebras can also be recast and generalized in the language of section 3. Since this is as straightforward to write out as it is lengthy and technical to justify in all its rigour, we omit further details here.
Definition 3.8 ([21, Definition 3.5]).
Two given cocycle representations $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ are called operator homotopic, if there exists a unitary $W\in\mathcal{U}_{0}(\mathscr{D}_{(\varphi,\mathbbmss{u})})$ such that $(\varphi,\mathbbmss{u})=\operatorname{Ad}(W)\circ(\psi,\mathbbmss{u})$.
Remark 3.9.
Let $\mathfrak{C}$ be an eligible set. Suppose that two given elements $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v})\in\mathfrak{C}$ are operator homotopic. Just as it was the case for ordinary equivariant $KK$-theory, it follows that the pair $\big((\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v})\big)$ is an $(\mathfrak{C};\alpha,\beta)$-Cuntz pair that is $\mathfrak{C}$-homotopic to the degenerate pair $\big((\varphi,\mathbbmss{u}),(\varphi,\mathbbmss{u})\big)$ by connecting the chosen unitary $W$ above to the unit. Hence $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v})]=0$ in $KK^{G}(\mathfrak{C};\alpha,\beta)$. This merely uses the fact that $\mathfrak{C}$ is closed under unitary equivalence.
In complete analogy to what was observed in [21, Section 3] as a Cuntz–Thomsen picture analog of Kasparov’s well-known observations about operator homotopy, we shall show in the rest of this section that the implication in section 3 has a stable converse. That is, if some equivariant $(\mathfrak{C};\alpha,\beta)$-Cuntz pair is $\mathfrak{C}$-homotopic to a degenerate one, then the two underlying cocycle representations are operator homotopic after adding a suitable representation in $\mathfrak{C}$.
Notation 3.10.
Assume that $\beta$ is a strongly stable action, witnessed by a sequence of isometries $r_{n}\in\mathcal{M}(B)^{\beta}$ with $\sum_{n=1}^{\infty}r_{n}r_{n}^{*}=\mathbf{1}$. Let us choose a family of matrix units $\{e_{k,\ell}\mid k,\ell\geq 1\}$ generating the $\mathrm{C}^{*}$-algebra $\mathcal{K}$ of compact operators on $\ell^{2}(\mathbb{N})$. Below, we denote by $\Lambda$ the equivariant isomorphism $\Lambda:(\mathcal{K}\otimes B,\operatorname{id}_{\mathcal{K}}\otimes\beta)\to(B,\beta)$ given by $\Lambda(e_{k,\ell}\otimes b)=r_{k}br_{\ell}^{*}$ for every $b\in B$ and $k,\ell\geq 1$.
Lemma 3.11.
Assume that $\mathfrak{C}$ is a downward closed set of cocycle representations from $(A,\alpha)$ to $(\mathcal{M}(B),\beta)$. Let $(\Omega,\mathbb{X})\in\mathfrak{C}[0,1]$. If $\theta:\mathcal{C}[0,1]\to\mathcal{M}(\mathcal{K})\cong\mathbb{B}(\ell^{2}(\mathbb{N}))$ is any faithful unital representation with $\theta(\mathcal{C}[0,1])\cap\mathcal{K}=\{0\}$, then the composition of
$(A,\alpha)\stackrel{{\scriptstyle(\Omega,\mathbb{X})}}{{\longrightarrow}}(\mathcal{M}(B[0,1]),\beta[0,1])\stackrel{{\scriptstyle\theta\otimes\operatorname{id}_{B}}}{{\longrightarrow}}(\mathcal{M}(\mathcal{K}\otimes B),\operatorname{id}_{\mathcal{K}}\otimes\beta)\stackrel{{\scriptstyle\Lambda}}{{\longrightarrow}}(\mathcal{M}(B),\beta)$
is a cocycle representation in $\mathfrak{C}$.
Proof.
Let $(\kappa,\mathbbmss{x})$ be the cocycle representation that arises from the composition above. Since $\mathfrak{C}$ is downward closed, it suffices to verify $(\kappa,\mathbbmss{x})\preccurlyeq\mathfrak{C}$.
We first claim that it suffices to prove this for a particular choice of $\theta$. Indeed, if $\theta_{1}$ and $\theta_{2}$ are two representations as in the statement, then denote by $(\kappa^{(j)},\mathbbmss{x}^{(j)})$ their associated compositions for $j=1,2$. It follows by Voiculescu’s theorem [47] that there exists a sequence of unitaries $U_{n}\in\mathcal{U}(\mathcal{M}(\mathcal{K}))$ such that $\operatorname{Ad}(U_{n})\circ\theta_{1}\to\theta_{2}$ in the point-norm topology. The sequence of unitaries $W_{n}=\Lambda(U_{n}\otimes\mathbf{1}_{\mathcal{M}(B)})\in\mathcal{M}(B)^{\beta}$ is then seen to satisfy the limit behavior
$\begin{array}[]{ccl}\operatorname{Ad}(W_{n})\circ\kappa^{(1)}&=&\Lambda\circ\operatorname{Ad}(U_{n}\otimes\mathbf{1}_{\mathcal{M}(B)})\circ(\theta_{1}\otimes\operatorname{id}_{B})\circ\Omega\\ &\to&\Lambda\circ(\theta_{2}\otimes\operatorname{id}_{B})\circ\Omega\ =\ \kappa^{(2)}\end{array}$
in the point-strict topology, as well as
$\begin{array}[]{ccl}W_{n}\mathbbmss{x}^{(1)}_{g}\beta_{g}(W_{n})^{*}&=&W_{n}\mathbbmss{x}_{g}W_{n}^{*}\\ &=&\Lambda(U_{n}\otimes\mathbf{1}_{\mathcal{M}(B)})\Lambda((\theta_{1}\otimes\operatorname{id}_{B})(\mathbb{X}_{g}))\Lambda(U_{n}^{*}\otimes\mathbf{1}_{\mathcal{M}(B)})\\ &\to&\Lambda((\theta_{2}\otimes\operatorname{id}_{B})(\mathbb{X}_{g}))\ =\ \mathbbmss{x}^{(2)}_{g}\end{array}$
strictly and uniformly over compact subsets of $G$. By [21, Proposition 4.7], it follows that $(\kappa^{(1)},\mathbbmss{x}^{(1)})$ is weakly contained in $(\kappa^{(2)},\mathbbmss{x}^{(2)})$. We conclude that if the claim holds for $\theta_{2}$ in place of $\theta$, then it also holds for $\theta_{1}$ as $\mathfrak{C}$ was downward closed. Since $\theta_{1}$ and $\theta_{2}$ were both arbitrary, we have verified that it suffices to prove the statement for a specific choice of $\theta$.
Let us choose a surjective map $\mathbb{N}\to\mathbb{Q}\cap[0,1]$, $n\mapsto t_{n}$. Let $(\delta_{n})_{n\geq 1}$ be the standard orthonormal basis for $\mathcal{H}=\ell^{2}(\mathbb{N})$ and define $\theta:\mathcal{C}[0,1]\to\mathcal{M}(\mathcal{K})=\mathcal{B}(\mathcal{H})$ via $\theta(f)(\xi)=\sum_{n=1}^{\infty}f(t_{n})\langle\delta_{n}\mid\xi\rangle\delta_{n}$ for all $f\in\mathcal{C}[0,1]$ and $\xi\in\ell^{2}(\mathbb{N})$. [^6] For this formula we adopt the convention that the inner product is linear in the second component. We proceed to prove the statement for this choice of $\theta$.
Let $\varepsilon>0$ be a constant, $b\in B$ a contraction, $\mathcal{F}\subset A$ a finite set and $K\subseteq G$ a compact subset. We may assume that $\mathcal{F}$ consists of contractions. We may choose a number $N$ and a positive contraction $e\in B$ such that $\|b(\mathbf{1}-b_{0})\|\leq\varepsilon/2$, where $b_{0}=\sum_{n=1}^{N}r_{n}er_{n}^{*}$. Let $P_{N}\in\mathcal{K}\subseteq\mathcal{M}(\mathcal{K}\otimes B)$ be given as $P_{N}=\sum_{n=1}^{N}e_{n,n}$. Then we have for all $f\in\mathcal{C}[0,1]$ that $P_{N}\theta(f)P_{N}=\sum_{n=1}^{N}f(t_{n})e_{n,n}$. Hence we have for all $a\in A$ that
$\begin{array}[]{ccl}b_{0}^{*}\kappa(a)b_{0}&=&\Lambda(P_{N}\otimes e)\kappa(a)\Lambda(P_{N}\otimes e)\\ &=&\Lambda\Big[(P_{N}\otimes e)(\theta\otimes\operatorname{id}_{B})(\Omega(a))(P_{N}\otimes e)\Big]\\ &=&\Lambda\Big[\sum_{n=1}^{N}e_{n,n}\otimes(e\Omega_{t_{n}}(a)e)\Big]\\ &=&\displaystyle\sum_{n=1}^{N}r_{n}e\Omega_{t_{n}}(a)er_{n}^{*}.\end{array}$
Furthermore we have for all $g\in G$ that
$\begin{array}[]{ccl}b_{0}^{*}\mathbbmss{x}_{g}\beta_{g}(b_{0})&=&b_{0}^{*}\cdot\Lambda((\theta\otimes\operatorname{id}_{B})(\mathbb{X}_{g}))\cdot\beta_{g}(b_{0})\\ &=&\Lambda\Big[(P_{N}\otimes e)(\theta\otimes\operatorname{id}_{B})(\mathbb{X}_{g})(P_{N}\otimes\beta_{g}(e))\Big]\\ &=&\Lambda\Big[\sum_{n=1}^{N}e_{n,n}\otimes(e\mathbb{X}_{g,t_{n}}\beta_{g}(e))\Big]\\ &=&\displaystyle\sum_{n=1}^{N}r_{n}e\mathbb{X}_{g,t_{n}}\beta_{g}(e)r_{n}^{*}.\end{array}$
If we consider $c_{j}=er_{n}^{*}b$ for all $j=1,\dots,N$, then we observe for all $a\in\mathcal{F}$ that
$b^{*}\kappa(a)b=_{\varepsilon}b^{*}b_{0}^{*}\kappa(a)b_{0}b=\sum_{n=1}^{N}c_{n}^{*}\Omega_{t_{n}}(a)c_{n}$
and for all $g\in K$ that
$b^{*}\mathbbmss{x}_{g}\beta_{g}(b)^{*}=_{\varepsilon}b^{*}b_{0}^{*}\mathbbmss{x}_{g}\beta_{g}(b_{0}b)=\sum_{n=1}^{N}c_{n}^{*}\mathbb{X}_{g,t_{n}}\beta_{g}(c_{n}).$
Since $(\Omega_{t},\mathbb{U}_{t})\in\mathfrak{C}$ for all $t\in[0,1]$ and the quadruple $(\varepsilon,b,\mathcal{F},K)$ was arbitrary, we get that indeed $(\kappa,\mathbbmss{x})\preccurlyeq\mathfrak{C}$. ∎
Lemma 3.12.
Let $\mathfrak{C}$ be a downward closed set of cocycle representations from $(A,\alpha)$ to $(\mathcal{M}(B),\beta)$. Suppose that $\big((\Phi,\mathbb{U}),(\Psi,\mathbb{V})\big)$ is a $(\mathfrak{C}[0,1];\alpha,\beta[0,1])$-Cuntz pair such that
$(\varphi,\mathbbmss{u}):=\operatorname{ev}_{0}\circ(\Phi,\mathbb{U})=\operatorname{ev}_{1}\circ(\Phi,\mathbb{U})=\operatorname{ev}_{0}\circ(\Psi,\mathbb{V}).$
Define $(\psi,\mathbbmss{v}):=\operatorname{ev}_{1}\circ(\Psi,\mathbb{V})$. Then there exists a cocycle representation $(\kappa,\mathbbmss{x})\in\mathfrak{C}$ such that $(\varphi,\mathbbmss{u})\oplus(\kappa,\mathbbmss{x})$ and $(\psi,\mathbbmss{v})\oplus(\kappa,\mathbbmss{x})$ are operator homotopic.
Proof.
We apply the construction and argument of [21, Notation 3.7, Lemma 3.8]. This gives a cocycle representation $(\kappa,\mathbbmss{x})$ such that $(\varphi,\mathbbmss{u})\oplus(\kappa,\mathbbmss{x})$ and $(\psi,\mathbbmss{v})\oplus(\kappa,\mathbbmss{x})$ are operator homotopic. Upon a close inspection of how $(\kappa,\mathbbmss{x})$ is constructed there [^7] This is admittedly quite lengthy to check when reading, but we omit the details here as this just comes from following the relevant steps outlined in [21] verbatim., this cocycle representation arises as a Cuntz sum of $(\varphi,\mathbbmss{u})$ and a composition of cocycle representations that is of the same form as in section 3. To be more precise, if one starts with a representation $\theta_{0}:\mathcal{C}[0,1]\to\mathcal{M}(\mathcal{K}(\mathcal{H}))$ satisfying the conclusion of [21, Theorem 2.15], then one sets
$\theta=\theta_{0}\oplus\theta_{0}:\mathcal{C}[0,1]\to\mathcal{M}(\mathcal{K}(\mathcal{H}\oplus\mathcal{H}))$
(as an ordinary direct sum) and
$(\Omega,\mathbb{X})=(\Psi,\mathbb{V})\oplus(\Phi,\mathbb{U}):(A,\alpha)\to(\mathcal{M}(B[0,1]),\beta[0,1])$
(as a Cuntz sum). Let $(\chi,\mathbbmss{y})$ be the cocycle representation that arises as a composition as in section 3 with these choices made. Evidently $(\Omega,\mathbb{X})\in\mathfrak{C}[0,1]$ and thus also $(\chi,\mathbbmss{y})\in\mathfrak{C}$. The construction and conclusion of [21, Notation 3.7, Lemma 3.8] then tells us that $(\kappa,\mathbbmss{x})=(\varphi,\mathbbmss{u})\oplus(\chi,\mathbbmss{y})\in\mathfrak{C}$ satisfies the desired property. ∎
We close this section by including a straightforward generalization of a known observation for $KK$-theory:
Proposition 3.13.
Let $\mathfrak{C}$ be an eligible set. Let $(\theta,\mathbbmss{y})\in\mathfrak{C}$ be an absorbing element. [^8] We remind that reader that if $B$ is separable, such an element always exists by [21, Theorem 4.16]. Then:
  1. (i)
    Every element $x\in\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)/{\sim_{h,\mathfrak{C}}}$ can be expressed as the equivalence class of a $(\mathfrak{C};\alpha,\beta)$-Cuntz pair $\big((\varphi,\mathbbmss{u}),(\theta,\mathbbmss{y})\big)$ for some absorbing element $(\varphi,\mathbbmss{u})\in\mathfrak{C}$.
  2. (ii)
    Every element $z\in KK^{G}(\mathfrak{C};\alpha,\beta)$ can be expressed as the equivalence class of an anchored $(\mathfrak{C};\alpha,\beta)$-Cuntz pair $\big((\varphi,\mathbbmss{u}),(\theta,\mathbbmss{y})\big)$ for some absorbing element $(\varphi,\mathbbmss{u})\in\mathfrak{C}$.
Proof.
This is the generalization of [21, Corollary 4.17]. We can copy the proof therein word for word to prove the statement here. The argument carries over because elements of $\mathfrak{C}$ are closed under unitary equivalence and Cuntz sums. ∎
4. $KK$-uniqueness
We remind the reader of the standing assumption stated earlier, which still applies to this and the subsequent sections: $G$ denotes an arbitrary second-countable, locally compact group, $\alpha:G\curvearrowright A$ is an action on a separable $\mathrm{C}^{*}$-algebra, and $\beta:G\curvearrowright B$ is a strongly stable action on a $\sigma$-unital $\mathrm{C}^{*}$-algebra.
Notation 4.1.
Let $(\varphi,\mathbbmss{u}):A\to\mathcal{M}(B)$ be a cocycle representation. We define
$\mathscr{C}_{(\varphi,\mathbbmss{u})}=\left\{x\in\mathcal{M}(B)\mid x\varphi(A)\cup\varphi(A)x\subseteq B,\left\{x-\beta^{\mathbbmss{u}}_{g}(x)\right\}_{g\in G}\subseteq B\right\}.$
When we compare with section 2, we see that $\mathscr{C}_{(\varphi,\mathbbmss{u})}\triangleleft\mathscr{D}_{(\varphi,\mathbbmss{u})}$ is an ideal. We may thus consider the quotient $\mathscr{B}_{(\varphi,\mathbbmss{u})}=\mathscr{D}_{(\varphi,\mathbbmss{u})}/\mathscr{C}_{(\varphi,\mathbbmss{u})}$.
Given an arbitrary *-homomorphism $\psi:A\to\mathcal{M}(B)$, we may view it as a cocycle representation with respect to $G=\{1\}$ and hence write $\mathscr{C}_{\psi}$ as short-hand for $\mathscr{C}_{(\psi,\mathbf{1})}$. Using this notation, we can also observe that when $(\varphi,\mathbbmss{u})$ is given as above, then $\mathscr{C}_{\varphi}$ is genuinely $\beta^{\mathbbmss{u}}$-invariant and $\mathscr{C}_{\varphi}\triangleleft\mathscr{D}^{w}_{(\varphi,\mathbbmss{u})}$ is an ideal.
The following lemma and many variations of it are well-known, but we include a proof for the reader’s convenience.
Lemma 4.2.
Let $E$ be any $\mathrm{C}^{*}$-algebra and $\theta:E\to E$ an endomorphism. Suppose that there exists an isometry $S\in\mathcal{M}(E)$ with $\theta(x)=SxS^{*}$ for all $x\in E$. Then one has for every ideal $I\triangleleft E$ that $\theta(I)\subseteq I$. For any inclusion of ideals $J\triangleleft I\triangleleft E$, denote by $\theta_{J}^{I}$ the induced endomorphism on $I/J$ via $\theta_{J}^{I}(x+J)=\theta(x)+J$ for all $x\in I$. It follows that $K_{*}(\theta^{I}_{J})=\operatorname{id}$ on $K_{*}(I/J)$.
Proof.
Given ideals $J\triangleleft I\triangleleft E$, it follows from the universal property of the multiplier algebra that there is a canonical *-homomorphism $\pi:\mathcal{M}(E)\to\mathcal{M}(I/J)$ given by $\pi(y)(b+J)=yb+J$ for all $b\in I$ and $y\in\mathcal{M}(E)$. Hence $\pi(S)\in\mathcal{M}(I)$ is an isometry implementing $\theta^{I}_{J}$. Therefore, it suffices to show the last part of the claim for $I=E$ and $J=0$.
Let $\{e_{k\ell}\}_{k,\ell=1,2}$ be the generating matrix units of the $2\times 2$ matrix algebra $M_{2}$. In order to prove the claim it suffices (by stability of $K$-theory) to show that the composition with the corner embedding $e_{11}\otimes\theta:E\to M_{2}(E)$ is homotopic to $e_{11}\otimes\operatorname{id}_{E}$. Choose some norm-continuous path of unitaries $W:[0,1]\to M_{2}\subseteq\mathcal{M}(M_{2}(E))$ such that $W_{0}=\mathbf{1}$ and $W_{1}=e_{21}+e_{12}$ (e.g., $W_{t}=\exp(\frac{\pi}{2}it(\mathbf{1}-W_{1}))$). The resulting continuous path of isometries $R_{t}=W_{t}(e_{11}\otimes S+e_{22}\otimes\mathbf{1})W_{t}^{*}$ satisfies $R_{0}=(e_{11}\otimes S+e_{22}\otimes\mathbf{1})$ and $R_{1}=(e_{11}\otimes\mathbf{1}+e_{22}\otimes S)$. The point-norm continuous path of embeddings $\theta_{t}:E\to M_{2}(E)$ for $t\in[0,1]$ given by $\theta_{t}(x)=R_{t}(e_{11}\otimes x)R_{t}^{*}$ then satisfies $\theta_{0}=e_{11}\otimes\theta$ and $\theta_{1}=e_{11}\otimes\operatorname{id}_{E}$. ∎
Lemma 4.3.
Suppose that $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ are two cocycle representations and that $(\varphi,\mathbbmss{u})$ weakly absorbs $(\psi,\mathbbmss{v})$. Consider the embedding
$\operatorname{id}\oplus 0:\mathscr{D}^{w}_{(\varphi,\mathbbmss{u})}\to\mathscr{D}^{w}_{(\varphi,\mathbbmss{u})\oplus(\psi,\mathbbmss{v})}\text{ given by }x\mapsto x\oplus 0\text{ (Cuntz addition).}$
Then $K_{*}(\operatorname{id}\oplus 0)$ is an isomorphism of abelian groups. Furthermore $\operatorname{id}\oplus 0$ descends to embeddings $\mathscr{D}_{(\varphi,\mathbbmss{u})}\to\mathscr{D}_{(\varphi,\mathbbmss{u})\oplus(\psi,\mathbbmss{v})}$ and $\mathscr{C}_{(\varphi,\mathbbmss{u})}\to\mathscr{C}_{(\varphi,\mathbbmss{u})\oplus(\psi,\mathbbmss{v})}$ that also induce isomorphisms in $K$-theory.
Proof.
Let $r_{1},r_{2}\in\mathcal{M}(B)^{\beta}$ be two isometries with $r_{1}r_{1}^{*}+r_{2}r_{2}^{*}=\mathbf{1}$. It shall be understood that Cuntz sums in this proof are formed with this pair. By assumption, there exists a unitary $U\in\mathcal{M}(B)$ such that $\operatorname{Ad}(U)\circ(\varphi\oplus\psi,\mathbbmss{u}\oplus\mathbbmss{v})$ agrees with $(\varphi,\mathbbmss{u})$ modulo $B$. The conjugation map $\kappa=\operatorname{Ad}(Ur_{1})$ then induces an embedding that is a composition
$\kappa:\mathscr{D}_{(\varphi,\mathbbmss{u})}\stackrel{{\scriptstyle\operatorname{Ad}(r_{1})}}{{\longrightarrow}}\mathscr{D}_{(\varphi\oplus\psi,\mathbbmss{u}\oplus\mathbbmss{v})}\stackrel{{\scriptstyle\operatorname{Ad}(U)}}{{\longrightarrow}}\mathscr{D}_{(\varphi,\mathbbmss{u})}.$
Since we have for all $a\in A$ that
$\varphi(a)Ur_{1}\equiv_{B}U(\varphi\oplus\psi)(a)r_{1}=Ur_{1}\varphi(a)$
and for all $g\in G$ that
$\beta_{g}^{\mathbbmss{u}}(Ur_{1})=\mathbbmss{u}_{g}\beta_{g}(U)r_{1}\mathbbmss{u}_{g}^{*}\equiv_{B}U(\mathbbmss{u}_{g}\oplus\mathbbmss{v}_{g})r_{1}\mathbbmss{u}_{g}^{*}=Ur_{1},$
it follows that $Ur_{1}\in\mathscr{D}_{(\varphi,\mathbbmss{u})}\subseteq\mathscr{D}_{(\varphi,\mathbbmss{u})}^{w}$. This element is obviously an isometry. It follows from section 4 that $\kappa$ induces isomorphisms on $K$-theory. Since the map $\operatorname{Ad}(U)$ above is a genuine isomorphism (section 2), it follows that $\operatorname{Ad}(r_{1})=\operatorname{id}\oplus 0:\mathscr{D}_{(\varphi,\mathbbmss{u})}\to\mathscr{D}_{(\varphi,\mathbbmss{u})\oplus(\psi,\mathbbmss{v})}$ induces an isomorphism on $K$-theory as well. It persists at the level of ideals such as the inclusion $\mathscr{C}_{(\varphi,\mathbbmss{u})}\to\mathscr{C}_{(\varphi,\mathbbmss{u})\oplus(\psi,\mathbbmss{v})}$. The same follows for $\mathscr{D}_{(\varphi,\mathbbmss{u})}^{w}$. ∎
Theorem 4.4.
Let $\alpha:G\curvearrowright A$ be an action on a separable $\mathrm{C}^{*}$-algebra and $\beta:G\curvearrowright B$ a strongly stable action on a $\sigma$-unital $\mathrm{C}^{*}$-algebra. Let $\mathfrak{C}$ be an eligible set of cocycle representations $(A,\alpha)\to(\mathcal{M}(B),\beta)$, and let $(\Phi,\mathbb{U})\in\mathfrak{C}$ be an absorbing element. Given any $U\in\mathcal{U}(\mathscr{D}_{(\Phi,\mathbb{U})})$, we consider the Cuntz pair $\xi(U)\in\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)$ given by $\operatorname{Ad}(U)\circ(\Phi,\mathbb{U})$ and $(\Phi,\mathbb{U})$. Then the assignment $\Xi:K_{1}(\mathscr{D}_{(\Phi,\mathbb{U})})\to\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)/_{\sim_{\mathfrak{C},h}}$ given by $\Xi([U])=\xi(U)$ defines an isomorphism of abelian groups.
Proof.
Let us first consider the map $\Omega:\mathcal{U}(\mathscr{D}_{(\Phi,\mathbb{U})})\to\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)/_{\sim_{\mathfrak{C},h}}$ given by $\Omega(U)=[\xi(U)]$. By section 2, $\mathscr{D}_{(\Phi,\mathbb{U})}$ is a properly infinite $K_{1}$-injective $\mathrm{C}^{*}$-algebra, hence the canonical map $\mathcal{U}(\mathscr{D}_{(\Phi,\mathbb{U})})/\mathcal{U}_{0}(\mathscr{D}_{(\Phi,\mathbb{U})})\to K_{1}(\mathscr{D}_{(\Phi,\mathbb{U})})$ is an isomorphism. Therefore it suffices to show that $\Omega$ is a well-defined surjective group homomorphism with kernel $\mathcal{U}_{0}(\mathscr{D}_{(\Phi,\mathbb{U})})$.
The fact that $\xi$ and therefore also $\Omega$ is well-defined is due to the definition of $\mathscr{D}_{(\Phi,\mathbb{U})}$. For any $U\in\mathcal{U}(\mathscr{D}_{(\Phi,\mathbb{U})})$, the cocycle representation $\operatorname{Ad}(U)\circ(\Phi,\mathbb{U})$ belongs to $\mathfrak{C}$ and is equal to $(\Phi,\mathbb{U})$ modulo $B$, thus indeed $\xi(U)\in\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)$. The fact that $\Omega$ is a group homomorphism follows from the following calculation for $U,V\in\mathcal{U}(\mathscr{D}_{(\Phi,\mathbb{U})})$:
$\begin{array}[]{ccl}\Omega(UV)&=&\big[\operatorname{Ad}(UV)\circ(\Phi,\mathbb{U}),(\Phi,\mathbb{U})\big]\\ &=&\big[\operatorname{Ad}(UV)\circ(\Phi,\mathbb{U}),\operatorname{Ad}(U)\circ(\Phi,\mathbb{U})\big]+\big[\operatorname{Ad}(U)\circ(\Phi,\mathbb{U}),(\Phi,\mathbb{U})\big]\\ &=&\big[\operatorname{Ad}(V)\circ(\Phi,\mathbb{U}),(\Phi,\mathbb{U})\big]+\big[\operatorname{Ad}(U)\circ(\Phi,\mathbb{U}),(\Phi,\mathbb{U})\big]\\ &=&\Omega(U)+\Omega(V).\end{array}$
For the surjectivity, let $\big((\psi,\mathbbmss{v}),(\theta,\mathbbmss{y})\big)\in\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)$ be a Cuntz pair. We shall employ an argument analogous to the proof of [21, Corollary 4.17]. Since $(\Phi,\mathbb{U})$ is an absorbing element of $\mathfrak{C}$, we have equivalences
$(\Phi,\mathbb{U})\oplus(\theta,\mathbbmss{y})\sim_{\mathrm{asymp}}(\Phi,\mathbb{U})\sim_{\mathrm{asymp}}(\Phi,\mathbb{U})\oplus(\psi,\mathbbmss{v}).$
Let $V:[0,\infty)\to\mathcal{U}(\mathcal{M}(B))$ be a norm-continuous path satisfying
  • $\Phi(a)=\displaystyle\lim_{t\to\infty}V_{t}(\Phi(a)\oplus\theta(a))V_{t}^{*}$ for all $a\in A$;
  • $\Phi(a)-V_{t}(\Phi(a)\oplus\theta(a))V_{t}^{*}\in B$ for all $a\in A$ and $t\geq 0$;
  • $\lim_{t\to\infty}\max_{g\in K}\ \|\mathbb{U}_{g}-U_{t}(\mathbb{U}_{g}\oplus\mathbbmss{y}_{g})\beta_{g}(U_{t})^{*}\|=0$ for all compact sets $K\subseteq G$;
  • $\mathbb{U}_{g}-U_{t}(\mathbb{U}_{g}\oplus\mathbbmss{x}_{g})\beta_{g}(U_{t})^{*}\in B$ for all $t\geq 0$ and $g\in G$.
From these properties we can read off that the degenerate $(\mathfrak{C};\alpha,\beta)$-Cuntz pair $\big((\Phi,\mathbb{U}),(\Phi,\mathbb{U})\big)$ is $\mathfrak{C}$-homotopic to $\big(\operatorname{Ad}(V_{0})\circ(\Phi\oplus\theta,\mathbb{U}\oplus\mathbbmss{y}),(\Phi,\mathbb{U})\big)$. Analogously, we find a norm-continuous path $W:[0,\infty)\to\mathcal{U}(\mathcal{M}(B))$ such that $\operatorname{Ad}(W_{t})\circ(\Phi\oplus\psi,\mathbb{U}\oplus\mathbbmss{v})$ converges to $(\Phi,\mathbb{U})$ in the same manner. Note that since $(\theta,\mathbbmss{y})$ and $(\psi,\mathbbmss{v})$ agree modulo $B$, the same is true for
$\begin{array}[]{cll}(\Phi,\mathbb{U})&\equiv_{B}&\operatorname{Ad}(V_{0})\circ(\Phi\oplus\theta,\mathbb{U}\oplus\mathbbmss{y})\\ &\equiv_{B}&\operatorname{Ad}(V_{0})\circ(\Phi\oplus\psi,\mathbb{U}\oplus\mathbbmss{v})\\ &\equiv_{B}&\operatorname{Ad}(V_{0}W_{0}^{*})\circ(\Phi,\mathbb{U}).\end{array}$
This means that the unitary $V_{0}W_{0}^{*}$ belongs to $\mathscr{D}_{(\Phi,\mathbb{U})}$. We compute in $\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)/_{\sim_{\mathfrak{C},h}}$ that
$\begin{array}[]{ccl}\big[(\psi,\mathbbmss{v}),(\theta,\mathbbmss{y})\big]&=&\big[(\Phi\oplus\psi,\mathbb{U}\oplus\mathbbmss{v}),(\Phi\oplus\theta,\Phi\oplus\mathbbmss{y})\big]\\ &=&\big[\operatorname{Ad}(W_{0}^{*})\circ(\Phi,\mathbb{U}),(\Phi\oplus\psi,\mathbb{U}\oplus\mathbbmss{v})\big]\\ &&+\big[(\Phi\oplus\psi,\mathbb{U}\oplus\mathbbmss{v}),(\Phi\oplus\theta,\Phi\oplus\mathbbmss{y})\big]\\ &&+\big[(\Phi\oplus\theta,\Phi\oplus\mathbbmss{y}),\operatorname{Ad}(V_{0}^{*})\circ(\Phi,\mathbb{U})\big]\\ &=&\big[\operatorname{Ad}(W_{0}^{*})\circ(\Phi,\mathbb{U}),\operatorname{Ad}(V_{0}^{*})\circ(\Phi,\mathbb{U})\big]\\ &=&\big[\operatorname{Ad}(V_{0}W_{0}^{*})\circ(\Phi,\mathbb{U}),(\Phi,\mathbb{U})\big]\ =\ \Omega(V_{0}W_{0}^{*}).\end{array}$
Since the pair $\big((\psi,\mathbbmss{v}),(\theta,\mathbbmss{y})\big)\in\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)$ was arbitrary, we see that $\Omega$ is surjective.
Lastly, let us determine the kernel of $\Omega$. By section 3 we have that $\mathcal{U}_{0}(\mathscr{D}_{(\Phi,\mathbb{U})})$ is contained in the kernel. We claim equality. Suppose that $U\in\mathcal{U}(\mathscr{D}_{(\Phi,\mathbb{U})})$ is given with $\Omega(U)=0$. This means that the $(\mathfrak{C};\alpha,\beta)$-Cuntz pair $\big(\operatorname{Ad}(U)\circ(\Phi,\mathbb{U}),(\Phi,\mathbb{U})\big)$ is $\mathfrak{C}$-homotopic to the degenerate pair $\big((\Phi,\mathbb{U}),(\Phi,\mathbb{U})\big)$. By section 3, it follows that there exists an element $(\kappa,\mathbbmss{x})\in\mathfrak{C}$ such that $(\Phi\oplus\kappa,\mathbb{U}\oplus\mathbbmss{x})$ and $\operatorname{Ad}(U\oplus\mathbf{1})\circ(\Phi\oplus\kappa,\mathbb{U}\oplus\mathbbmss{x})$ are operator homotopic. Without loss of generality, by adding more to $(\kappa,\mathbbmss{x})$ if necessary, we may assume that $(\Phi\oplus\kappa,\mathbb{U}\oplus\mathbbmss{x})$ is an infinite repeat. There exists $W\in\mathcal{U}_{0}(\mathscr{D}_{(\Phi\oplus\kappa,\mathbb{U}\oplus\mathbbmss{x})})$ such that
$\operatorname{Ad}(W(U\oplus\mathbf{1}))\circ(\Phi\oplus\kappa,\mathbb{U}\oplus\mathbbmss{x})=(\Phi\oplus\kappa,\mathbb{U}\oplus\mathbbmss{x}).$
In other words, we have that $W(U\oplus\mathbf{1})\in\big(\mathcal{M}(B)\cap(\Phi\oplus\kappa)(A)^{\prime}\big)^{\beta^{\mathbb{U}\oplus\mathbbmss{x}}}$. By subsection 1.1, the $K_{1}$-class of $W(U\oplus\mathbf{1})$ in $\mathscr{D}_{(\Phi\oplus\kappa,\mathbb{U}\oplus\mathbbmss{x})}$ is therefore trivial. As $W$ is homotopic to the unit, we have $U\oplus\mathbf{1}]=0$ in $K_{1}(\mathscr{D}_{(\Phi\oplus\kappa,\mathbb{U}\oplus\mathbbmss{x})})$. By section 4, we must have $[U]=0$ in $K_{1}(\mathscr{D}_{(\Phi,\mathbb{U})})$ and hence $U\in\mathcal{U}_{0}(\mathscr{D}_{(\Phi,\mathbb{U})})$ with section 2. This finishes the proof. ∎
Before we state our main result, we state an important technical lemma that we will use in its proof.
Lemma 4.5 (see [21, Corollary 5.4]).
Let $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ be two cocycle representations. Suppose that there exists a norm-continuous path of unitaries $U:[0,\infty)\to\mathcal{U}(\mathscr{D}_{(\varphi,\mathbbmss{u})})$ with $U_{0}=\mathbf{1}$ and
$\lim_{t\to\infty}\Big(\|\psi(a)-U_{t}\varphi(a)U_{t}^{*}\|+\max_{g\in K}\|\mathbbmss{v}_{g}-U_{t}\mathbbmss{u}_{g}\beta_{g}(U_{t})^{*}\|\Big)=0$
for all $a\in A$ and every compact set $K\subseteq G$. Then $(\varphi,\mathbbmss{u})$ and $(\psi,\mathbbmss{v})$ are strongly asymptotically unitarily equivalent.
This brings us to the main result of the article, which is a generalized dynamical $KK$-uniqueness theorem that improves upon the dynamical stable uniqueness theorem [21, Theorem 6.4]; see [12, Theorems 3.8, 3.10] and [6, Theorem 5.15] for the non-dynamical case. It also entails the uniqueness theorem for nuclear and/or ideal-related $KK$-theory as treated in [18, Theorem 13.9]. The content of this section also represents a somewhat more self-contained approach to answer [6, Question 5.17] in the positive, as well as an assortment of conceivable variants of that question for other versions of $KK$-theory. Restricting its statement to the case of separable $\mathrm{C}^{*}$-algebras, one obtains Introduction.
Theorem 4.6.
Let $\alpha:G\curvearrowright A$ be an action on a separable $\mathrm{C}^{*}$-algebra and $\beta:G\curvearrowright B$ a strongly stable action on a $\sigma$-unital $\mathrm{C}^{*}$-algebra. Let $\mathfrak{C}$ be an eligible set of cocycle representations $(A,\alpha)\to(\mathcal{M}(B),\beta)$. Let $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v})\in\mathfrak{C}$ be two absorbing elements that form an anchored $(\mathfrak{C};\alpha,\beta)$-Cuntz pair. Then $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v})]=0$ in $KK^{G}(\mathfrak{C};\alpha,\beta)$ if and only if $(\varphi,\mathbbmss{u})$ and $(\psi,\mathbbmss{v})$ are strongly asymptotically unitarily equivalent.
Proof.
Since the “if” part is trivial, we only need to prove the “only if” part. Assume $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v})]=0$ in $KK^{G}(\mathfrak{C};\alpha,\beta)$. Since the pair was anchored, this is equivalent to $(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v})]=0$ in the group $\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)/_{\sim_{\mathfrak{C},h}}$. Since $(\varphi,\mathbbmss{u})$ and $(\psi,\mathbbmss{v})$ were both absorbing elements of $\mathfrak{C}$, we have
$(\varphi,\mathbbmss{u})\sim_{\mathrm{asymp}}(\varphi,\mathbbmss{u})\oplus(\psi,\mathbbmss{v})\sim_{\mathrm{asymp}}(\psi,\mathbbmss{v}).$
Choose a norm-continuous path $U:[1,\infty)\to\mathcal{U}(\mathcal{M}(B))$ satisfying
  1. $\bullet$
    $\varphi(a)=\displaystyle\lim_{t\to\infty}U_{t}\psi(a)U_{t}^{*}$ for all $a\in A$;
  2. $\bullet$
    $\varphi(a)-U_{t}\psi(a)U_{t}^{*}\in B$ for all $a\in A$ and $t\geq 1$;
  3. $\bullet$
    $\lim_{t\to\infty}\max_{g\in K}\ \|\mathbbmss{u}_{g}-U_{t}\mathbbmss{v}_{g}\beta_{g}(U_{t})^{*}\|=0$ for all compact sets $K\subseteq G$;
  4. $\bullet$
    $\mathbbmss{u}_{g}-U_{t}\mathbbmss{v}_{g}\beta_{g}(U_{t})^{*}\in B$ for all $t\geq 1$ and $g\in G$.
Given that $(\varphi,\mathbbmss{u})$ and $(\psi,\mathbbmss{v})$ agree modulo $B$, the second and the fourth items imply that $U_{t}\in\mathscr{D}_{(\varphi,\mathbbmss{u})}$ for all $t\geq 1$. This path induces the equality $0=[(\varphi,\mathbbmss{u}),(\psi,\mathbbmss{v})]=[\operatorname{Ad}(U_{1})\circ(\psi,\mathbbmss{v}),(\psi,\mathbbmss{v})]$. In light of section 4, this means $0=[U_{1}]$ in $K_{1}(\mathscr{D}_{(\varphi,\mathbbmss{u})})$, which by section 2 means that $U_{1}\in\mathcal{U}_{0}(\mathscr{D}_{(\Phi,\mathbb{U})})$. We may thus extend the given path $(U_{t})_{t\geq 1}$ to a norm-continuous path $U:[0,\infty)\to\mathcal{U}(\mathscr{D}_{(\varphi,\mathbbmss{u})})$ with $U_{0}=\mathbf{1}$. This path satisfies the conditions required by section 4 and hence it follows that $(\varphi,\mathbbmss{u})$ and $(\psi,\mathbbmss{v})$ are strongly asymptotically unitarily equivalent. ∎
5. A generalized Paschke-type duality
This section is devoted to proving a generalized Paschke-type duality that unifies earlier results of a similar nature. We start with some preparatory observations.
Lemma 5.1.
Let $(\varphi,\mathbbmss{u}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ be a cocycle representation and $x\in\mathscr{D}^{w}_{(\varphi,\mathbbmss{u})}$. Then for every $a\in A$, the map $g\mapsto\varphi(a)\beta_{g}^{\mathbbmss{u}}(x)$ is norm-continuous.
Proof.
Let $h\in G$ and $a\in A$. We then have $\lim_{g\to h}\alpha_{g}(\alpha_{h}^{-1}(a))=a$ in norm and therefore
$\lim_{g\to h}\varphi(a)\beta_{g}^{\mathbbmss{u}}(x)=\lim_{g\to h}\beta_{g}^{\mathbbmss{u}}(\varphi(\alpha_{h}^{-1}(a))x).$
Since $a$ and $h$ were arbitrary and $\alpha$ is point-norm continuous on $A$, we see that the claim is equivalent to saying that the assignment $g\mapsto\beta^{\mathbbmss{u}}_{g}(\varphi(a)x)$ is norm-continuous for every $a\in A$. Let $\mathcal{Q}(B)$ denote the corona algebra of $B$, let $\bar{\beta}^{\mathbbmss{u}}$ be the algebraic $G$-action induced by $\beta^{\mathbbmss{u}}$, let $\bar{\varphi}:A\to\mathcal{Q}(B)$ be the induced homomorphism, and write $\bar{x}=x+B\in\mathcal{Q}(B)$. By [4, Theorem 2.1], it suffices to show that the map $g\mapsto\bar{\beta}^{\mathbbmss{u}}_{g}(\bar{\varphi}(a)\bar{x})$ is norm-continuous as a map into $\mathcal{Q}(B)$. The condition $x,\varphi(A)]\subseteq B$ means $\bar{x}\in\mathcal{Q}(B)\cap\bar{\varphi}(A)^{\prime}$. The condition $\varphi(A)(\beta^{\mathbbmss{u}}_{g}(x)-x)\subseteq B$ for all $g\in G$ means that the element
$x^{\sharp}=\bar{x}+(\mathcal{Q}(B)\cap\bar{\varphi}(A)^{\perp})\in(\mathcal{Q}(B)\cap\bar{\varphi}(A)^{\prime})/(\mathcal{Q}(B)\cap\bar{\varphi}(A)^{\perp})$
is in the fixed point algebra of the action induced by $\bar{\beta}^{\mathbbmss{u}}$, which we shall denote by $E$. We have an induced (algebraic) $G$-equivariant *-homomorphism $\kappa:(A\otimes_{\max}E,\alpha\otimes\operatorname{id})\to(\mathcal{Q}(B),\bar{\beta}^{\mathbbmss{u}})$ given by
$\kappa\big(a\otimes\big(y+(\mathcal{Q}(B)\cap\bar{\varphi}(A)^{\perp})\big)\big)=\bar{\varphi}(a)y$
for all $a\in A$ and $y\in\mathcal{Q}(B)\cap\bar{\varphi}(A)^{\prime}$ with $y+(\mathcal{Q}(B)\cap\bar{\varphi}(A)^{\perp})\in E$. By equivariance of $\kappa$, it follows for all $g,h\in G$ that
$\begin{array}[]{ccl}\|\bar{\beta}_{g}^{\mathbbmss{u}}(\bar{\varphi}(a)\bar{x})-\bar{\beta}_{h}^{\mathbbmss{u}}(\bar{\varphi}(a)\bar{x})\|&=&\|\kappa\big((\alpha_{g}(a)-\alpha_{h}(a))\otimes x^{\sharp}\big)\|\\ &\leq&\|(\alpha_{g}(a)-\alpha_{h}(a))\otimes x^{\sharp}\|\\ &\leq&\|x\|\|\alpha_{g}(a)-\alpha_{h}(a)\|.\end{array}$
This shows the desired continuity property and finishes the proof. ∎
Proposition 5.2.
Let $(\varphi,\mathbbmss{u}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ be a cocycle representation. Keeping in mind section 2 and section 4, one has $\mathscr{C}_{(\varphi,\mathbbmss{u})}=\mathscr{D}_{(\varphi,\mathbbmss{u})}\cap\mathscr{C}_{\varphi}$. The resulting inclusion of quotients $\mathscr{B}_{(\varphi,\mathbbmss{u})}=\mathscr{D}_{(\varphi,\mathbbmss{u})}/\mathscr{C}_{(\varphi,\mathbbmss{u})}\subseteq\mathscr{D}^{w}_{(\varphi,\mathbbmss{u})}/\mathscr{C}_{\varphi}$ is an isomorphism.
Proof.
The only thing to show is that this map is surjective. Concretely, we have to show that for every $x\in\mathscr{D}^{w}_{(\varphi,\mathbbmss{u})}$, there exists $y\in\mathscr{D}_{(\varphi,\mathbbmss{u})}$ such that $x-y\in\mathscr{C}_{\varphi}$. Given $x$, we consider the strictly continuous map $\zeta:G\to\mathcal{M}(B)$ given by $\zeta(g)=\beta^{\mathbbmss{u}}_{g}(x)-x$. By definition of $\mathscr{D}^{w}_{(\varphi,\mathbbmss{u})}$, we have $\zeta(G)\varphi(A)\subseteq B$. Furthermore, every map of the form $\zeta(\_)\varphi(a)$ is norm-continuous by section 5. We apply Kasparov’s technical theorem [31, Theorem 1.4] for $B$ in place of $J$ with $G$-action $\beta^{\mathbbmss{u}}$, $\varphi(A)$ in place of $A_{1}$, $\zeta$ in place of $\varphi$, and $A_{2}=\Delta=0$. This yields two positive contractions $M,N\in\mathcal{M}^{\beta^{\mathbbmss{u}}}(B)$ with $M+N=\mathbf{1}$ satisfying $\zeta(G)N\subset B$ and $\varphi(A)M\subset B$. Clearly $M,N\in\mathscr{D}_{(\varphi,\mathbbmss{u})}$. The element $y=xN\in\mathscr{D}_{\varphi}$ then satisfies for all $g\in G$ that
$\beta^{\mathbbmss{u}}_{g}(y)-y\equiv_{B}(\beta^{\mathbbmss{u}}_{g}(x)-x)N=\zeta(g)N\in B.$
Hence $y\in\mathscr{D}_{(\varphi,\mathbbmss{u})}$. On the other hand, we have for all $a\in A$ that
$(x-y)\varphi(a)=x(1-N)\varphi(a)=xM\varphi(a)\in B.$
In other words, we see that indeed $x-y\in\mathscr{C}_{\varphi}$, which finishes the proof. ∎
The main idea behind the next lemma and its corollary is analogous to a similar argument in the proof of [45, Lemma 6.4]. We note that the assumption below, which we impose on a cocycle representation, is an analog of (but not identical to) what Thomsen therein refers to as a “saturated” equivariant *-homomorphism.
Lemma 5.3.
Let $(\varphi,\mathbbmss{u}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ be a cocycle representation that weakly absorbs $(0,\mathbbmss{u})$ and contains itself at infinity. Then the inclusion $\mathscr{C}_{(\varphi,\mathbbmss{u})}\subseteq\mathcal{M}^{\beta^{\mathbbmss{u}}}(B)$ induces an embedding in $K$-theory with a split.
Proof.
When comparing definitions, notice that $\mathscr{C}_{(0,\mathbbmss{u})}=\mathcal{M}^{\beta^{\mathbbmss{u}}}(B)$. Let $r_{1},r_{2}\in\mathcal{M}(B)^{\beta}$ be two isometries with $r_{1}r_{1}^{*}+r_{2}r_{2}^{*}=\mathbf{1}$, which we use to form Cuntz sums. By assumption, there exists a unitary $U\in\mathcal{M}(B)$ such that $\operatorname{Ad}(U)\circ(\varphi\oplus 0,\mathbbmss{u}\oplus\mathbbmss{u})$ agrees with $(\varphi,\mathbbmss{u})$ modulo $B$. Consider the norm-continuous path of isometries given by
$r_{1+t}:=\sqrt{1-t}\cdot r_{1}+\sqrt{t}\cdot r_{2},\quad t\in[0,1].$
Let us consider the point-norm continuous family of embeddings $\kappa_{t}:\mathscr{C}_{(\varphi,\mathbbmss{u})}\to\mathcal{M}(B)$ for $t\in[0,1]$ given by $\kappa_{t}(x)=r_{1+t}xr_{1+t}^{*}$. Let $t\in[0,1]$ and $x\in\mathscr{C}_{(\varphi,\mathbbmss{u})}$. We notice for all $a\in A$ that
$\kappa_{t}(x)(\varphi(a)\oplus 0)=r_{1+t}x\cdot\sqrt{1-t}\cdot\varphi(a)r_{1}^{*}\in B.$
Futhermore, we have for all $g\in G$ that
[…]
In other words, for each $t\in[0,1]$ the map $\kappa_{t}$ takes values in $\mathscr{C}_{(\varphi\oplus 0,\mathbbmss{u}\oplus\mathbbmss{u})}$. Let $\iota:\mathscr{C}_{(\varphi,\mathbbmss{u})}\to\mathcal{M}^{\beta^{\mathbbmss{u}}}(B)$ be the inclusion map. Then the path $\{\kappa_{t}\}_{t\in[0,1]}$ witnesses the fact that the diagram
${\mathscr{C}_{(\varphi,\mathbbmss{u})}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\iota}$$\scriptstyle{\operatorname{id}\oplus 0}$${\mathcal{M}^{\beta^{\mathbbmss{u}}}(B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{0\oplus\operatorname{id}}$${\mathscr{C}_{(\varphi\oplus 0,\mathbbmss{u}\oplus\mathbbmss{u})}}$
commutes up to homotopy. The arrow $\operatorname{id}\oplus 0$ induces an isomorphism in $K$-theory by section 4. We can thus see that $K_{*}(\operatorname{id}\oplus 0)^{-1}\circ K_{*}(0\oplus\operatorname{id})$ is a left inverse of $K_{*}(\iota)$. This finishes the proof. ∎
Corollary 5.4.
Let $\psi:A\to B$ be any *-homomorphism that contains itself at infinity and absorbs the zero representation. Then $\mathscr{C}_{\psi}$ has vanishing $K$-theory. In particular, if $(\varphi,\mathbbmss{u}):(A,\alpha)\to(\mathcal{M}(B),\beta)$ is any cocycle representation such that $\varphi$ contains itself at infinity and absorbs the zero representation, then the quotient map $\mathscr{D}_{(\varphi,\mathbbmss{u})}^{w}\to\mathscr{B}_{(\varphi,\mathbbmss{u})}$ (resulting from section 5) induces an isomorphism in $K$-theory.
Proof.
Upon considering $G=\{1\}$, we may view $\psi$ as a cocycle representation in a trivial way. The assumptions on $\psi$ translate into the assumptions required by section 5. Hence the inclusion $\mathscr{C}_{\psi}\subseteq\mathcal{M}(B)$ is injective in $K$-theory. The $K$-groups of $\mathcal{M}(B)$ are known to vanish; see [7, Lemma 3.2]. This can also be deduced from applying subsection 1.1 to the special case $(\varphi,\mathbbmss{u})=(0,\mathbf{1})$. The “in particular” part is an immediate consequence of the general properties of the $K$-theory functor. ∎
Below follows our version of the Paschke duality theorem. We note that for $G=\{1\}$, most of the terms simplify and we recover Thomsen’s [44, Theorem 3.2]. His dynamical generalization [45, Theorem 6.2] is very similar, but was phrased in the language of genuinely equivariant representations, whereas the result below is tailored to fit within the framework of cocycle representations.
Theorem 5.5.
Let $\alpha:G\curvearrowright A$ be an action on a separable $\mathrm{C}^{*}$-algebra and $\beta:G\curvearrowright B$ a strongly stable action on a $\sigma$-unital $\mathrm{C}^{*}$-algebra. Let $\mathfrak{C}$ be an eligible set of cocycle representations $(A,\alpha)\to(\mathcal{M}(B),\beta)$, and let $(\Phi,\mathbb{U})\in\mathfrak{C}$ be an absorbing element. Consider the isomorphism $\Xi:K_{1}(\mathscr{D}_{(\Phi,\mathbb{U})})\to\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)/_{\sim_{\mathfrak{C},h}}$ from section 4. Then $\Xi$ restricts to an isomorphism $K_{1}(\mathscr{C}_{(\Phi,\mathbb{U})})\cong H_{\beta}$. This induces isomorphisms $K_{1}(\mathscr{D}_{(\Phi,\mathbb{U})}^{w})\cong K_{1}(\mathscr{B}_{(\Phi,\mathbb{U})})\cong KK^{G}(\mathfrak{C};\alpha,\beta)$ such that the following diagram commutes:
${H_{\beta}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)/_{\sim_{\mathfrak{C},h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathfrak{f}}$${KK^{G}(\mathfrak{C};\alpha,\beta)}$${K_{1}(\mathscr{C}_{(\Phi,\mathbb{U})})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${K_{1}(\mathscr{D}_{(\Phi,\mathbb{U})})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Xi}\scriptstyle{\cong}$${K_{1}(\mathscr{B}_{(\Phi,\mathbb{U})})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${K_{1}(\mathcal{M}^{\beta^{\mathbb{U}}}(B))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$$\scriptstyle{\Xi_{0}}$${K_{1}(\mathscr{D}_{(\Phi,\mathbb{U})}^{w})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$
Proof.
Before we begin with the proof, we point out a basic fact that we use below for convenience. All of the $\mathrm{C}^{*}$-algebras appearing in the diagram above are either unital properly infinite $\mathrm{C}^{*}$-algebras or ideals in such a $\mathrm{C}^{*}$-algebra, hence any element in the $K_{1}$-group has a representative in the unitary group of the $\mathrm{C}^{*}$-algebra. This makes it possible to avoid higher matrix amplifications below.
We begin by first explaining in what sense one should understand the leftmost upward arrow $\Xi_{0}$. One has $\mathcal{M}^{\beta^{\mathbb{U}}}(B)=\mathscr{D}_{(0,\mathbb{U})}$ upon comparing definitions. Since $(\Phi,\mathbb{U})$ is an absorbing element of $\mathfrak{C}$ and $\mathfrak{C}$ contains all elements of the form $(0,\mathbb{V})$, it follows that $(\Phi,\mathbb{U})\sim_{\mathrm{asymp}}(\Phi\oplus 0,\mathbb{U}\oplus\mathbb{V})$ for every $\beta$-cocycle $\mathbb{V}$. Since this implies $(0,\mathbb{U})\sim_{\mathrm{asymp}}(0,\mathbb{U}\oplus\mathbb{V})$, it follows that $(0,\mathbb{U})$ is an absorbing cocycle representation when viewed as one with domain $(A,\alpha)=(\{0\},\operatorname{id}_{\{0\}})$. Thus we may apply section 4 in this sense and conclude that
$\Xi_{0}:K_{1}(\mathcal{M}^{\beta^{\mathbb{U}}}(B))\to H_{\beta}=\mathbb{E}^{G}(\operatorname{id}_{\{0\}},\beta)/_{\sim_{h}}$
given by $W]\mapsto\big[\operatorname{Ad}(W)\circ(0,\mathbb{U}),(0,\mathbb{U})\big]$ is an isomorphism.
On the other hand, when we take $U]\in K_{1}(\mathscr{D}_{(\Phi,\mathbb{U})})$, then we see
$H_{\beta}\ni(\mathfrak{f}\circ\Xi)([U])=\mathfrak{f}\big[\operatorname{Ad}(U)\circ(\Phi,\mathbb{U}),(\Phi,\mathbb{U})\big]=\big[\operatorname{Ad}(U)\circ(0,\mathbb{U}),(0,\mathbb{U})\big].$
This means that the composition $\mathfrak{f}\circ\Xi$ agrees with the composition of $K_{1}(\mathscr{D}_{(\Phi,\mathbb{U})})\to K_{1}(\mathcal{M}^{\beta^{\mathbb{U}}}(B))\stackrel{{\scriptstyle\Xi_{0}}}{{\to}}H_{\beta}$.
Now let $W]\in K_{1}(\mathscr{C}_{(\Phi,\mathbb{U})})$ for some $W\in\mathcal{U}(\mathbf{1}+\mathscr{C}_{(\Phi,\mathbb{U})})$. Let $r_{1},r_{2}\in\mathcal{M}(B)^{\beta}$ be two isometries with $r_{1}r_{1}^{*}+r_{2}r_{2}^{*}=\mathbf{1}$, which we use to form Cuntz sums. Exactly as in the proof of section 5, we consider the norm-continuous path of isometries $r_{1+t}:=\sqrt{1-t}\cdot r_{1}+\sqrt{t}\cdot r_{2}$ for $t\in[0,1]$ and the point-norm continuous family of embeddings $\kappa_{t}:\mathscr{C}_{(\Phi,\mathbb{U})}\to\mathscr{C}_{(\Phi\oplus 0,\mathbb{U}\oplus\mathbb{U})}$ for $t\in[0,1]$ given by $\kappa_{t}(x)=r_{1+t}xr_{1+t}^{*}$.
The assignment $0,1]\ni t\mapsto r_{t}Wr_{t}^{*}+(\mathbf{1}-r_{t}r_{t}^{*})$ induces a homotopy between $W\oplus\mathbf{1}$ and $\mathbf{1}\oplus W$ inside $\mathcal{U}(\mathbf{1}+\mathscr{C}_{(\Phi\oplus 0,\mathbb{U}\oplus\mathbb{U})})$. This yields a chain of $\mathfrak{C}$-homotopies
[…]
Hence the composition of the maps $K_{1}(\mathscr{C}_{(\Phi,\mathbb{U})})\to K_{1}(\mathcal{M}^{\beta^{\mathbb{U}}}(B))\stackrel{{\scriptstyle\Xi}}{{\to}}H_{\beta}\to\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)/_{\sim_{\mathfrak{C},h}}$ is equal to the composition of the maps $K_{1}(\mathscr{C}_{(\Phi,\mathbb{U})})\to K_{1}(\mathscr{D}_{(\Phi,\mathbb{U})})\stackrel{{\scriptstyle\Xi}}{{\to}}\mathbb{E}^{G}(\mathfrak{C};\alpha,\beta)/_{\sim_{\mathfrak{C},h}}$. This induces the homomorphism
$K_{1}(\mathscr{C}_{(\Phi,\mathbb{U})})\to H_{\beta}$
such that the whole left half of the diagram commutes. It is injective by section 5. We claim that it is also surjective. Given any $x\in H_{\beta}$, we may appeal to the surjectivity of $\Xi_{0}$ and find a unitary $V\in\mathcal{M}^{\beta^{\mathbb{U}}}$ such that $x=\big[\operatorname{Ad}(V)\circ(0,\mathbb{U}),(0,\mathbb{U})\big]$. Since the map
$K_{1}(\operatorname{id}\oplus 0):K_{1}(\mathscr{C}_{(\Phi,\mathbb{U})})\to K_{1}(\mathscr{C}_{(\Phi\oplus 0,\mathbb{U}\oplus\mathbb{U})})$
is an isomorphism by section 4, we find a unitary $W\in\mathcal{U}(\mathbf{1}+\mathscr{C}_{(\Phi,\mathbb{U})})$ with $W\oplus\mathbf{1}]=[\mathbf{1}\oplus V]$ in $K_{1}(\mathscr{C}_{(\Phi\oplus 0,\mathbb{U}\oplus\mathbb{U})})$. By the $K_{1}$-injectivity of $\mathscr{D}_{(\Phi\oplus 0,\mathbb{U}\oplus\mathbb{U})}$, we have that $W\oplus\mathbf{1}$ and $\mathbf{1}\oplus V$ are homotopic inside the unitary group of $\mathscr{D}_{(\Phi\oplus 0,\mathbb{U}\oplus\mathbb{U})}$. The analogous computation to the above then shows that
$\big[\operatorname{Ad}(W)\circ(\Phi,\mathbb{U}),(0,\mathbb{U})\big]=\big[\operatorname{Ad}(V)\circ(0,\mathbb{U}),(0,\mathbb{U})\big]=x.$
This takes care of the left half of the diagram.
The isomorphism $K_{1}(\mathscr{B}_{(\Phi,\mathbb{U})})\to KK^{G}(\mathfrak{C};\alpha,\beta)$ making the rest of the diagram commute is then uniquely induced because
$K_{1}(\mathscr{C}_{(\Phi,\mathbb{U})})\hookrightarrow K_{1}(\mathscr{D}_{(\Phi,\mathbb{U})})\twoheadrightarrow K_{1}(\mathscr{B}_{(\Phi,\mathbb{U})})$
is also a short exact sequence due to $\mathfrak{f}$ inducing a left inverse for the first arrow. We have already observed in section 5 and section 5 that the arrow $K_{1}(\mathscr{D}^{w}_{(\Phi,\mathbb{U})})\to K_{1}(\mathscr{B}_{(\Phi,\mathbb{U})})$ is a well-defined isomorphism. This finishes the proof. ∎
Remark 5.6.
Let $\mathfrak{C}$ be an eligible set of cocycle representations $(A,\alpha)\to(\mathcal{M}(B),\beta)$ and let $(\Phi,\mathbb{U})\in\mathfrak{C}$ be an absorbing element. The isomorphism $K_{1}(\mathscr{B}_{(\Phi,\mathbb{U})})\cong KK^{G}(\mathfrak{C};\alpha,\beta)$ from section 5 is, strictly speaking, only the first of two possible versions of Paschke duality in our context. It can be viewed as a generalization of [45, Theorem 6.2] in the language of cocycle representations instead of genuinely equivariant ones. The second version of Paschke duality should relate $K_{0}(\mathscr{B}_{(\Phi,\mathbb{U})})$ to a generalized version of the group $KK^{G}_{1}(\alpha,\beta)$ and/or $\operatorname{Ext}^{G}(\alpha,\beta)$ built from cocycle representations or Busby maps related to $\mathfrak{C}$. A rigorous investigation of how such generalized equivariant Ext- or $KK_{1}$-groups would actually look like is beyond the scope of this article, but the apparent template for this kind of second Paschke duality would be [45, Theorems 6.3, 6.5]; see also [44, Section 3]. For $\mathfrak{C}$ being the set of all cocycle representations, the related constructions are available in the literature, and such a result could in this case perhaps be achieved with some extra work if one appeals to [43, Lemma 4.4]. Since that result and its proof appeal to various technical results in the Kasparov picture of $KK^{G}_{1}$ (including the general form of the Kasparov product and its properties from [30]), it is not immediately clear how to generalize the argument to the present context in a self-contained manner.
Acknowledgements
The author was supported by the European Research Council under the European Union’s Horizon Europe research and innovation programme (ERC grant AMEN–101124789). For the purpose of open access, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.
I would like to thank Stuart White, Shanshan Hua, James Gabe, Sergio Girón Pacheco and Robert Neagu for fruitful conversations regarding the topic of this work, and for providing feedback on earlier versions of the article. This work has benefited from participation in the programme Topological groupoids and their $\mathrm{C}^{*}$-algebras in July 2025 and I thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for the support and hospitality. That programme was supported by EPSRC grant EP/V521929/1.